### v3

# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

Abstract
While many advances in time series models focus exclusively on numerical data, research on multimodal time series, particularly those involving contextual textual information, remains in its infancy. With recent progress in large language models and time series learning, we revisit the integration of paired texts with time series through the Platonic Representation Hypothesis (Huh et al., 2024), which posits that representations of different modalities converge to shared spaces. In this context, we identify that time-series-paired texts may naturally exhibit periodic properties that closely mirror those of the original time series. Building on this insight, we propose a novel framework, Texts as Time Series (TaTS), which considers the time-series-paired texts to be auxiliary variables of the time series. TaTS can be plugged into any existing numerical-only time series models and effectively enable them to handle time series data with paired texts. Through extensive experiments on both multimodal time series forecasting and imputation tasks across benchmark datasets with various existing time series models, we demonstrate that TaTS can enhance multimodal predictive performance without modifying model architectures. Our Code is available at https://github.com/iDEA-iSAIL-Lab-UIUC/TaTS.
1 Introduction
Figure 1: Mean Square Error of modeling frameworks of time series with paired texts. Full results in Appendix E.2.
Time series modeling plays an important role in a wide range of real-world applications, including finance (Sezer et al., 2020), healthcare (Zhang et al., 2024b), climate (Fu et al., 2024), and energy systems (Kotzur et al., 2017; Li et al., 2015). While extensive research has focused on approaches that rely solely on the numerical values of time series (Zhou et al., 2021; Liu et al., 2024c; Wu et al., 2023b; Wang et al., 2024c), real-world scenarios often involve additional modalities that co-occur with the time series and can provide valuable complementary information (De Baets and Harvey, 2023; Rai et al., 2023; Kyei and Antwi, 2017; Wang et al., 2024d).
In scenarios like pandemic policymaking, economic planning, or investment strategies, textual information can provide explanations, updates, or external factors that influence the underlying numerical patterns. However, research on effectively leveraging data from other modalities paired with time series remains in its early stages. In this work, we focus on time-series paired with texts at each timestamp, a common data format where textual descriptions are associated with time series at each timestamp in a parallel manner, as illustrated in Figure 3 (left). For instance, during a pandemic, infection rates are often accompanied by government announcements and news reports in real-time (Cinelli et al., 2020). On the one hand, numerical-only models overlook valuable contextual information that may influence or explain the patterns in the time series. On the other hand, the current state-of-the-art approach (Liu et al., 2024a) disregards the unique positional characteristics that time-series-paired texts may inherently possess. Such limitations raise a pivotal question:
What unique attributes may characterize time-series-paired texts, and how can they be systematically integrated to improve time series modeling and predictions?
In this paper, motivated by the Platonic Representation Hypothesis (PRH) (Huh et al., 2024), we pioneer the exploration of effectively leveraging paired texts to enrich time series analysis. We identify an intriguing phenomenon, which we term “Chronological Textual Resonance” (CTR): depending on data quality, time-series-paired texts may exhibit periodic patterns that closely reflect the temporal dynamics of their corresponding numerical time series. More specifically, despite variations in expressions, the hidden representations of two periodicity-lagged texts associated with time series may demonstrate high similarities, revealing a deeper alignment between textual and numerical modalities. We attribute this phenomenon to the fact that the paired texts inherently evolve in response to the dynamics of the time series itself. Further, we introduce TT-Wasserstein, a new metric designed to measure the level of CTR and quantify alignment quality.
Building on these insights, we propose Texts as Time Series (TaTS), a simple yet effective framework for integrating paired texts to enhance multimodal time series modeling. Previous studies have shown that different variables in one multivariate time series exhibit similar periodicity properties (Zhang and Yan, 2023; Wang et al., 2024d; Yi et al., 2024), and CTR suggests that time-series-paired texts follow a similar pattern. This observation implies that paired texts can be considered as special auxiliary variables to augment the original time series. Motivated by this, TaTS first transforms the paired textual information into a lower-dimensional representation, then combines the original time series with the textual representations as new variables to form an augmented time series. This augmented time series is subsequently fed into existing time series models, allowing them to capture both numerical and textual dynamics. TaTS offers two key benefits: (i) it effectively captures the evolving positional characteristics of texts paired with a time series; and (ii) it functions as a plug-in module, maintaining compatibility with existing time series models. Empirically, the proposed TaTS achieves state-of-the-art performance on both forecasting and imputation tasks. Notably, we observe that a higher CTR level (i.e., a lower TT-Wasserstein) correlates with greater improvements compared to numerical-only modeling. In summary, our contributions are:
  • We revisit multimodal time series with the PRH and uncover a previously overlooked phenomenon, termed Chronological Textual Resonance (CTR), that time-series-paired texts may exhibit periodic patterns closely aligned with their corresponding numerical time series. Further, we propose TT-Wasserstein to quantify the level of CTR and the alignment quality.
  • Based on this phenomenon, we propose a plug-and-play multimodal time series forecasting framework, Texts as Time Series (TaTS), which transforms text representations into auxiliary variables, seamlessly integrating them into existing time series models.
  • Experiments on diverse benchmark datasets and multiple time series models demonstrate TaTS’s superior performance without requiring modifications to model architectures.
2 Preliminary
We use calligraphic letters (e.g., $\mathcal{A}$) for sets and bold capital letters for matrices (e.g., $\mathbf{A}$). For matrix indices, $\mathbf{A}[i,j]$ denotes the entry in the $i^{\textrm{th}}$ row and the $j^{\textrm{th}}$ column. For a vector $\mathbf{v}$, $v[i:j]$ represents the sub-vector sliced from the $i^{\textrm{th}}$ to the $j^{\textrm{th}}$ position, inclusively. $\mathbf{A}[i,:]$ returns the $i^{th}$ row in $\mathbf{A}$ and $\mathbf{A}[:i]$ returns the first $i$ rows of $\mathbf{A}$. In this paper, we focus on both forecasting and imputation.
Time Series Forecasting. A time series is denoted as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}\in\mathbb{R}^{T\times N}$, where $T$ represents the number of time steps and $N$ denotes the number of variables. $\mathbf{x}_{i}$ is the time series sequence of the $i^{\textrm{th}}$ variable. When $N>1$, the time series is referred to as a multivariate time series. Let $\mathbf{X}_{a:b}$ represent the time slice of the series from timestamp $a$ to $b$, i.e., $\mathbf{X}_{a:b}=\{\mathbf{x}_{1}[a:b],\mathbf{x}_{2}[a:b],\dots,\mathbf{x}_{N}[a:b]\}$. The task of time series forecasting is to predict the future $H$ steps:
$\widehat{\mathbf{X}}_{T+1:T+H}=\mathcal{F}\left(\mathbf{X}_{1:T};\theta_{\text{forecast}}\right)\in\mathbb{R}^{H\times N},$ (1)
where $\mathcal{F}$ denotes the mapping function, and $\theta_{\text{forecast}}$ denotes the learnable parameters of $\mathcal{F}$.
Time Series Imputation. The goal of imputation is to estimate missing values in the time series $\mathbf{X}$, where the missing entries are denoted by a binary mask $\mathbf{M}\in\{0,1\}^{T\times N}$. Specifically, $\mathbf{M}_{t,n}=1$ indicates $\mathbf{X}_{t,n}$ being observed, $\mathbf{M}_{t,n}=0$ indicates $\mathbf{X}_{t,n}$ being missing. Formally,
$\widehat{\mathbf{X}}^{\text{Imputed}}=\mathcal{G}\left(\mathbf{X}\odot\mathbf{M},\mathbf{M};\theta_{\text{impute}}\right)\in\mathbb{R}^{T\times N},$ (2)
where $\widehat{\mathbf{X}}^{\text{Imputed}}$ represents the imputed time series, $\mathcal{G}$ denotes the imputation function, $\theta_{\text{impute}}$ denotes its learnable parameters, and $\odot$ represents the element-wise multiplication. The imputation process aims to recover the missing entries such that $\widehat{\mathbf{X}}^{\text{Imputed}}\approx\mathbf{X}$ with respect to the actual $\mathbf{X}$ values.
Extending Time Series with Paired Texts. In addition to the numerical time series $\mathbf{X}\in\mathbb{R}^{T\times N}$, the dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$ includes textual information $\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}$, where each $s_{t}$ represents the text at timestamp $t$. Each $s_{t}$ is a string that can be tokenized into a sequence of tokens, i.e., $\text{Tokenize}(s_{t})=\{w_{t,1},w_{t,2},\dots,w_{t,L_{t}}\}$, where $L_{t}$ denotes the number of tokens in the text at time $t$. The textual data can be transformed into numerical representations using a textual encoder $\mathcal{H}_{\text{text}}$:
$\mathbf{e}_{t}=\mathcal{H}_{\text{text}}(s_{t};\theta_{\text{text}})\in\mathbb{R}^{d_{\text{text}}},$ (3)
where $\mathbf{e}_{t}$ is the encoded text representation at time $t$, $d_{\text{text}}$ is text embedding dimension, and $\theta_{\text{text}}$ are encoder parameters. In this work, we leverage pre-trained large language models to encode the texts.
(a) Economy
(b) Social Good
(c) Traffic
Figure 2: By overlaying the top frequencies of paired texts (vertical dashed lines) onto the amplitude spectrum of the time series, it is observed that the time-series-paired texts exhibit similar periodic properties that closely mirror those of the original time series. We term this phenomenon Chronological Textual Resonance. More Details are provided in Appendix A.
3 Chronological Textual Resonance (CTR)
The association of time series at each timestamp may imbue time-series-paired texts with unique characteristics that can be effectively harnessed through an appropriate design.
The Platonic Representation Hypothesis. PRH (Huh et al., 2024) posits that different modalities describing the same object converge towards a shared, latent representation. Extending this hypothesis, if time series and paired text both describe the same changing event, their representations are dynamic projections from a common underlying source, and should exhibit similar periodic properties.
To illustrate this hypothesis for time series with paired texts, we analyze three real-world datasets (Liu et al., 2024a), including (i) Economy: The time series represents trade data of the U.S., while the texts describe the general economic conditions of the country. (ii) Social Good: The time series captures the unemployment rate in the U.S., and the texts include detailed unemployment reports. (iii) Traffic: The time series reflects monthly travel volume trends from the U.S. Department of Transportation, with corresponding texts derived from traffic volume reports issued by the same department.
For each dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, we employ Fourier Transform (Nussbaumer, 1982; Sneddon, 1995) to analyze the frequency components of time series data and identify its dominant periodic components, illustrated by the blue curves in Figure 2. To examine the periodicity of texts, we embed each $s_{t}\in\mathbf{S}$ to obtain the text embedding $e_{t}$ at timestamp $t$. Then, we compute their lag-similarity, defined as $d_{l}=\sum_{t}\mathrm{cos}(e_{t},e_{t+L})$ where $L$ is the lag and $\mathrm{cos}(\cdot,\cdot)$ represents the cosine similarity. If the text embeddings exhibit a significant periodic pattern, the lag-similarity $d_{l}$ will also fluctuate periodically as the lag $l$ increases (proof in Proposition A.1). Finally, we identify major frequencies (with the largest amplitudes) of the texts by applying FFT to text lag-similarity, and mark them with red dashed lines, as shown in Figure 2. Detailed process is provided in Appendix A. We find that the major frequencies of the paired texts closely match those of the time series. Specifically, the paired texts also show periodicity of 12 (frequency 0.083) for monthly sampled time series, indicating that the paired texts exhibit periodicity that is strongly aligned with the temporal dynamics of the time series.
Figure 3: Texts as Time Series (TaTS) framework. As paired texts may exhibit behaviors similar to accompanying variables in a time series, TaTS transforms the paired texts into auxiliary variables. These variables augment the numerical sequence, forming a unified multimodal sequence that can be seamlessly integrated into any existing time series model.
Why may CTR happen? We present three key reasons for the observed alignment in periodicity between time series and their paired texts: (i) Shared External Drivers: Both the time series and their paired texts are often influenced by common external factors, such as seasonal changes, recurring events, or societal and economic cycles. These shared drivers naturally induce periodicity in both modalities. (ii) Influence of Time Series on Texts: Paired texts often serve as contextual reflections of the underlying time series, adapting and evolving in response to numerical trends. For instance, news articles or government reports accompanying economic indicators are frequently updated in response to the numerical trends. (iii) Texts Contain Additional Variables with Aligned Periodicity: Paired texts often include additional variables that are closely related to the time series. For example, if the time series represents economic GDP data, the accompanying texts may reference related variables such as stock market indices or inflation rates. These related variables often exhibit periodicity patterns aligned with the time series and affect the periodicity of the paired texts.
Table 1: TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets and their random shuffle.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
Original 0.026 0.025 0.022 0.049 0.027 0.035 0.307 0.233 0.302
TS shuffled 0.088 0.032 0.098 0.054 0.069 0.102 0.320 0.268 0.358
Text shuffled 0.106 0.037 0.099 0.053 0.072 0.104 0.312 0.277 0.364
Quantifying CTR level. Of course we cannot expect every time series with paired texts at each timestamp to exhibit periodicity similarity. In some cases, texts aligned with timestamps may not provide meaningful or complementary information, for example, daily lottery winning numbers. To this end, we propose a new metric for CTR, TT-Wasserstein, defined as the Wasserstein distance between the normalized spectra of time series and texts. Formally, based on Wasserstein distance $W(P,Q)$ (Kantorovich, 1960), for time series with paired texts dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, after computing normalized frequencies and amplitudes $\tilde{f}_{\text{texts}},\tilde{f}_{\text{ts}},\tilde{a}_{\text{texts}},\tilde{a}_{\text{ts}}$,
$\begin{split}\text{TT-Wasserstein}(\mathcal{D})&=W(P_{\text{texts}},P_{\text{ts}})=\inf_{\gamma\in\Pi(P_{\text{text}},P_{\text{ts}})}\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}\cdot\left|\tilde{f}_{\text{texts}}^{(i)}-\tilde{f}_{\text{ts}}^{(j)}\right|\\ \text{Subject to}&\quad\gamma_{ij}\geq 0,\quad\sum_{j=1}^{m}\gamma_{ij}=\tilde{a}_{\text{texts}}^{(i)},\quad\sum_{i=1}^{n}\gamma_{ij}=\tilde{a}_{\text{ts}}^{(j)}\end{split}$ (4)
Intuitively, $\text{TT-Wasserstein}(\mathcal{D})$ quantifies the discrepancy between the spectral distributions of paired texts and time series data. By design, a lower value of $\text{TT-Wasserstein}(\mathcal{D})$ should indicate a higher alignment between the textual and numerical modalities. To validate this, we compute our TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets, whose web-retrieved texts are filtered, disentangled, and summarized for relevance and alignment. We also compute TT-Wasserstein on time-series-shuffled or text-shuffled versions of the datasets to disrupt alignment. As shown in Table 1, the corrupted Time-MMD datasets yield much larger Wasserstein distances compared to the original datasets. This finding suggests that TT-Wasserstein can serve as an indicator for the alignment between modalities and a gauge of dataset quality. In Section 5, we will show that TT-Wasserstein can even predict the potential effectiveness of our proposed framework. Additionally, because TT-Wasserstein is an empirical statistical metric, we further analyze its sensitivity and estimation stability in Appendix A.1.
4 Texts as Time Series
An overview of the proposed Texts as Time Series (TaTS) is illustrated in Figure 3.
Concurrent Texts are Secretly Auxiliary Variables. As we elucidated in Section 3, the properties of concurrent texts closely align with those of variables in a multivariate time series: similar to numerical variables, concurrent texts are influenced by shared external drivers and interact dynamically with the time series. Furthermore, mapping concurrent texts to structured variables enables capturing hidden variables embedded within the concurrent texts.
Given the dataset $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}\}$, TaTS first embed the texts using a text encoder $\mathcal{H}_{\text{text}}$ to obtain text embeddings $\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\in\mathbb{R}^{d_{\text{text}\times T}}$. Since the text embedding dimension $d_{\text{text}}$ is typically much larger than the number of variables in the time series, we reduce the dimensionality of the text embeddings by applying a Multi-Layer Perceptron (MLP), mapping them into a lower-dimensional space of reduced dimensionality $d_{\text{mapped}}$.
$\mathbf{z}_{t}=\text{MLP}(\mathbf{e}_{t};\theta_{\text{MLP}})\in\mathbb{R}^{d_{\text{mapped}}}$ (5)
Unifying by Plugging-in a Time Series Model. The resulting mapped embeddings $\mathbf{Z}=\{\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{T}\}\in\mathbb{R}^{d_{\text{mapped}\times T}}$ are then treated as auxiliary variables in the time series. Specifically, $\mathbf{Z}$ is concatenated with $\mathbf{X}$ to form a unified multimodal sequence:
$\mathbf{U}=[\mathbf{X};\mathbf{Z}^{\intercal}]_{\text{dim=1}}\in\mathbb{R}^{T\times(N+d_{\text{mapped}})}$ (6)
The unified sequence $\mathbf{U}$ is then passed into an existing time series model for downstream tasks. Here, we formulate the example of forecasting the next $H$ steps of the time series
$\widehat{\mathbf{X}}_{T+1:T+H}=\mathcal{F}\left(\mathbf{U}_{1:T};\theta_{\text{forecast}}\right)[:N]\in\mathbb{R}^{H\times N}$ (7)
where $\mathcal{F}(\cdot;\theta_{\text{forecast}})$ denotes the time series forecasting model with parameters $\theta_{\text{forecast}}$, and $[:N]$ extracts the first $N$ variables corresponding to the original time series.
Finally, we joint train the time series model $\theta_{\text{forecast}}$ as well as the mapping MLP $\theta_{\text{MLP}}$ using the Mean Squared Error (MSE) loss.
$\mathcal{L}_{\text{forecast}}(\mathbf{X},\widehat{\mathbf{X}})=\frac{1}{H\cdot N}\sum_{t=T+1}^{T+H}\sum_{i=1}^{N}\left(\mathbf{X}_{t,i}-\widehat{\mathbf{X}}_{t,i}\right)^{2}$ (8)
where $\mathbf{X}_{t,i}$ and $\widehat{\mathbf{X}}_{t,i}$ represent the ground truth and predicted values of the $i^{\textrm{th}}$ variable at time step $t$, respectively. TaTS for imputation follows a similar process and is omitted here due to space constraints. Detailed algorithms for both forecasting and imputation are provided in Appendix C.
To conclude this section, we note that while some existing literatures make similar but limited attempts to model covariates in multivariate time series or multiple time series. Our TaTS framework differentiates from those approaches by unifying multi-modality using deep learning architectures. In Section 5, we compare TaTS with these existing approaches.
5 Experiment
Table 2: Time-series forecasting performance. Compared to numerical-only unimodal modeling and $M$-TSFLib, TaTS consistently and significantly enhances existing time series models to effectively handle paired texts. The results are averaged across all prediction lengths. Full results reported in Appendix E.3. Promotion (positive or negative) denotes the percentage reduction (or increase) in MSE or MAE achieved by TaTS compared to the best-performing baseline.
Models iTransformer PatchTST Crossformer DLinear FEDformer FiLM Autoformer Informer Transformer
(2024c) (2023) (2023) (2023) (2022b) (2022a) (2021) (2021) (2017)
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Agriculture
Uni-modal
0.122 0.251 0.120 0.247 0.323 0.406 0.223 0.354 0.138 0.286 0.139 0.256 0.158 0.297 0.599 0.630 0.354 0.434
$M$-TSFLib
0.112 0.230 0.114 0.233 0.218 0.313 0.218 0.355 0.131 0.275 0.140 0.258 0.158 0.288 0.313 0.414 0.249 0.352
TaTS (ours)
0.109 0.229 0.114 0.235 0.212 0.312 0.214 0.351 0.131 0.276 0.135 0.251 0.125 0.266 0.255 0.348 0.191 0.313
Promotion
[…]
Climate
Uni-modal
1.183 0.871 1.220 0.895 1.124 0.837 1.190 0.872 1.192 0.893 1.270 0.911 1.131 0.865 1.110 0.841 1.092 0.839
$M$-TSFLib
1.044 0.810 1.030 0.806 1.002 0.772 1.104 0.837 1.011 0.797 1.179 0.871 1.053 0.827 1.001 0.792 0.998 0.783
TaTS (ours)
1.028 0.804 1.004 0.798 0.938 0.755 0.931 0.759 0.926 0.760 0.945 0.772 0.980 0.789 0.930 0.756 0.920 0.753
Promotion
[…]
Economy
Uni-modal
0.014 0.096 0.017 0.105 0.758 0.828 0.058 0.192 0.042 0.166 0.025 0.129 0.071 0.207 1.325 1.110 0.584 0.711
$M$-TSFLib
0.011 0.086 0.014 0.096 0.250 0.458 0.058 0.192 0.035 0.153 0.026 0.129 0.058 0.192 0.432 0.618 0.213 0.416
TaTS (ours)
0.008 0.077 0.009 0.079 0.219 0.419 0.021 0.117 0.015 0.101 0.009 0.080 0.024 0.121 0.299 0.512 0.079 0.232
Promotion
[…]
Energy
Uni-modal
0.269 0.375 0.269 0.376 0.293 0.406 0.291 0.396 0.240 0.351 0.278 0.385 0.319 0.428 0.309 0.425 0.297 0.405
$M$-TSFLib
0.267 0.378 0.272 0.379 0.291 0.407 0.289 0.395 0.238 0.354 0.279 0.385 0.320 0.428 0.301 0.413 0.293 0.405
TaTS (ours)
0.265 0.376 0.258 0.371 0.279 0.394 0.283 0.388 0.237 0.355 0.271 0.379 0.314 0.430 0.284 0.396 0.279 0.395
Promotion
[…]
Environment
Uni-modal
0.441 0.494 0.552 0.537 0.551 0.581 0.558 0.591 0.503 0.549 0.577 0.543 0.599 0.598 0.459 0.512 0.460 0.511
$M$-TSFLib
0.421 0.478 0.459 0.501 0.427 0.488 0.429 0.502 0.423 0.486 0.478 0.490 0.452 0.502 0.424 0.480 0.425 0.481
TaTS (ours)
0.267 0.369 0.273 0.371 0.284 0.403 0.298 0.428 0.275 0.378 0.272 0.371 0.285 0.387 0.285 0.406 0.276 0.397
Promotion
[…]
Health
Uni-modal
1.587 0.817 1.652 0.855 1.535 0.827 1.737 0.848 1.486 0.909 1.982 1.005 1.962 1.039 1.278 0.773 1.378 0.776
$M$-TSFLib
1.446 0.816 1.347 0.797 1.273 0.744 1.541 0.800 1.252 0.792 1.675 0.949 1.494 0.887 1.215 0.740 1.218 0.748
TaTS (ours)
1.315 0.744 1.283 0.753 1.226 0.728 1.412 0.787 1.244 0.791 1.421 0.838 1.409 0.861 1.183 0.752 1.142 0.718
Promotion
[…]
Security
Uni-modal
115.94 5.660 112.85 5.371 126.96 6.277 109.11 4.711 114.48 5.158 115.55 5.487 115.27 5.118 131.78 6.623 131.35 6.582
$M$-TSFLib
116.34 5.532 112.84 5.369 125.72 6.183 108.03 4.712 113.73 5.107 109.19 4.897 111.44 4.973 128.95 6.415 128.47 6.378
TaTS (ours)
112.05 5.151 109.69 5.019 125.16 6.148 107.92 4.676 107.37 4.718 107.85 4.736 108.49 4.787 126.66 6.276 124.58 6.134
Promotion
[…]
Social Good
Uni-modal
1.212 0.483 1.097 0.495 0.865 0.467 1.151 0.712 0.979 0.476 1.261 0.654 1.278 0.701 0.870 0.504 0.910 0.484
$M$-TSFLib
1.197 0.520 1.073 0.515 0.837 0.398 1.083 0.673 0.962 0.462 1.236 0.626 1.229 0.670 0.839 0.457 0.856 0.461
TaTS (ours)
0.987 0.452 0.972 0.465 0.779 0.412 1.006 0.622 0.888 0.430 1.104 0.626 1.195 0.666 0.810 0.459 0.807 0.419
Promotion
[…]
Traffic
Uni-modal
0.213 0.238 0.188 0.242 0.214 0.376 0.230 0.359 0.205 0.264 0.215 0.314 0.212 0.298 0.202 0.355 0.209 0.346
$M$-TSFLib
0.199 0.347 0.178 0.230 0.188 0.334 0.209 0.330 0.193 0.238 0.207 0.300 0.212 0.272 0.172 0.299 0.171 0.295
TaTS (ours)
0.187 0.217 0.172 0.209 0.168 0.286 0.188 0.300 0.173 0.212 0.176 0.248 0.177 0.229 0.164 0.281 0.164 0.274
Promotion
[…]
In this section, we empirically validate the effectiveness of the proposed TaTS framework. In particular, we use GPT2 (Radford et al., 2019) encoder to embed the paired texts. We also validate the performance of TaTS with other text encoders across different datasets in section 5.2 and Appendix E.7. Implementation details are provided in Appendix D.4.
Datasets. We evaluate our framework on 18 real-world datasets from Time-MMD (Liu et al., 2024a), FNSPID Dong et al. (2024), and FNF Wang et al. (2024b). The datasets span diverse domains with sample frequencies ranging from daily to weekly and monthly. More details are in Appendix D.1.
Time Series Models and Baselines. To demonstrate the compatibility of TaTS with existing time series models, we integrate TaTS with 9 widely used models across different categories, including (i) Transformer-based models: iTransformer (Liu et al., 2024c), PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021) and Transformer (Vaswani et al., 2017). (ii) Linear models: DLinear (Zeng et al., 2023). (iii) Frequency-based models: FEDformer (Zhou et al., 2022b), FiLM (Zhou et al., 2022a).
We compare our TaTS framework with a wide array of baselines, including (a) numerical-only uni-modal modeling, which ignores the paired texts and utilizes only the numerical time series with the given time series model; (b) MM-TSFLib  (Liu et al., 2024a), a recently library proposed together with Time-MMD datasets for multimodal time series forecasting; (c) covariate-based methods: N-BEATS (Oreshkin et al., 2020), N-HiTS (Challu et al., 2022); (d) convolution-based method TCN (Bai et al., 2018); (e) a recent multimodal time series foundation model ChatTime (Wang et al., 2025).
Metrics. We evaluate the performance of multimodal time series modeling using MSE, MAE, RMSE, MAPE, and MSPE. Due to space constraints, we report only the MSE and MAE results in the main paper, while the results for the other metrics are provided in Appendix E.
5.1 Main Results
Improved Forecasting over Uni-modal Modeling and MM-TSFLib. Table 2 presents the performance results for the time series forecasting task on Time-MMD datasets. For each dataset and time series model, we report the average performance across four different prediction lengths, and the full results for each prediction length are provided in Appendix E. For datasets with relatively few samples, we perform short-term forecasting with prediction lengths of $\{6,8,10,12\}$. In contrast, for datasets with a larger number of samples, we perform long-term forecasting with prediction lengths of $\{48,96,192,336\}$. From the results, compared to uni-modal modeling or MM-TSFLib, our TaTS consistently achieves the best performance across all datasets. Notably, by plugging in TaTS to various existing time series models, it achieves an average performance improvement of over $5\%$ on 6 out of 9 datasets and delivers a remarkable performance boost of over $30\%$ on the largest dataset, Environment. The results also demonstrate that TaTS is highly compatible with a wide range of existing time series forecasting models, consistently delivering performance improvements across all of them in both long-term forecasting and short-term forecasting tasks. We provide a visualization of the performance boost in Appendix E.2, showcasing the improvements achieved by different time series models on each dataset.
Table 3: Imputation task performance. Full results in Appendix E.4. Promotion: percentage reduction by TaTS over the best-performing baseline.
Models PatchTST DLinear FiLM
(2023) (2023) (2022a)
Metric
MSE
MAE
MSE
MAE
MSE
MAE
Climate
Uni-modal
1.111 0.846 0.969 0.801 1.123 0.829
$M$-TSFLib
1.010 0.821 0.963 0.802 1.130 0.833
TaTS (ours)
0.878 0.720 0.912 0.757 0.820 0.718
Promotion
13.1% 12.3% 5.3% 5.5% 27.0% 13.4%
Economy
Uni-modal
0.029 0.138 0.057 0.190 0.077 0.209
$M$-TSFLib
0.026 0.137 0.061 0.196 0.075 0.203
TaTS (ours)
0.017 0.045 0.045 0.171 0.054 0.168
Promotion
34.6% 67.2% 21.0% 11.2% 28.0% 17.2%
Traffic
Uni-modal
0.210 0.339 0.245 0.417 0.175 0.311
$M$-TSFLib
0.189 0.341 0.179 0.335 0.169 0.288
TaTS (ours)
0.131 0.248 0.134 0.297 0.137 0.242
Promotion
30.7% 26.8% 25.1% 11.3% 18.9% 16.0%
Improved Time Series Imputation. We evaluate the performance of our TaTS framework on the imputation task using the Climate, Economy, and Traffic datasets, each with an imputation length of 24. Though MM-TSFLib only supports forecasting tasks, we extend it to serve as an imputation baseline by applying a similar linear interpolation. We select one representative time series model from each category and present the results in Table 3. The results demonstrate that TaTS consistently enhances the imputation capabilities of existing time series models, achieving improvements of up to $30\%$ compared to baseline methods.
From the above results, effectively leveraging the text modality provides significant benefits when paired texts are available, and our TaTS framework achieves notable improvements over baseline methods that disregard positional information in time-series-paired texts.
Table 4: TaTS compared with several baselines on a variety of datasets. Best results are bolded and second-best results are underlined. Full results in Table 28.
Methods TaTS (ours) TaTS (ours) TaTS (ours) N-BEATS N-HiTS TCN ChatTime GPT4MTS
+ iTransformer + PatchTST + FiLM (2020) (2022) (2018) (2025) (2024)
Datasets
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Time-MMD: Multimodal Time Series (2024a) Agriculture 0.109 0.229 0.114 0.235 0.135 0.251 3.267 1.458 1.852 1.032 4.168 1.797 0.508 0.447 0.327 0.393
Climate 1.028 0.804 1.004 0.798 0.945 0.772 1.093 0.861 1.103 0.858 1.098 0.866 1.568 1.019 1.127 0.873
Economy 0.008 0.077 0.009 0.079 0.009 0.080 1.010 0.920 0.444 0.584 5.546 2.349 0.049 0.166 0.014 0.096
Energy 0.265 0.376 0.258 0.371 0.271 0.379 0.329 0.424 0.372 0.463 0.430 0.512 0.305 0.417 0.269 0.378
Environment 0.267 0.369 0.273 0.371 0.272 0.371 0.518 0.573 0.522 0.583 0.854 0.713 0.580 0.594 0.348 0.422
Health 1.315 0.744 1.283 0.753 1.421 0.838 1.660 0.938 1.666 0.898 1.938 0.970 1.668 0.909 1.766 0.875
Security 112.054 5.151 109.693 5.019 107.850 4.736 130.065 6.618 138.124 6.978 136.596 6.873 133.106 6.887 119.407 5.198
Social Good 0.987 0.452 0.972 0.465 1.104 0.626 1.316 0.665 1.272 0.649 1.314 0.961 1.264 0.652 1.414 0.559
Traffic 0.187 0.217 0.172 0.209 0.176 0.248 0.347 0.464 0.268 0.382 0.708 0.744 0.363 0.429 0.195 0.246
FNSPID: Company Stock Price (2024) Delta Airlines (DAL) 0.087 0.197 0.086 0.192 0.095 0.201 0.286 0.444 0.226 0.379 0.294 0.466 0.098 0.202 0.093 0.201
IBM (IBM) 0.564 0.501 0.550 0.490 0.891 0.695 1.105 0.777 1.215 0.806 1.936 1.123 0.602 0.536 0.639 0.525
JPMorgan (JPM) 1.693 0.970 1.872 0.990 2.513 1.096 2.419 1.175 3.426 1.232 3.764 1.711 2.037 1.043 2.133 1.122
NVIDIA (NVDA) 0.043 0.141 0.048 0.156 0.050 0.174 0.272 0.434 0.122 0.262 0.457 0.574 0.053 0.161 0.054 0.159
Pfizer (PFE) 0.326 0.416 0.347 0.422 0.448 0.477 0.676 0.572 0.824 0.663 0.628 0.559 0.408 0.477 0.369 0.439
Tesla (TSLA) 0.142 0.281 0.158 0.297 0.110 0.244 0.188 0.332 0.254 0.416 3.476 1.851 0.158 0.300 0.181 0.303
FNF: Forecast with News (2024b) Bitcoin Price 2.609 1.112 2.339 1.045 2.775 1.142 141.738 10.355 159.019 10.535 57.826 6.744 3.590 1.344 2.721 1.152
Web Traffic 17.744 2.712 17.964 2.653 18.110 2.772 22.233 3.017 23.263 3.011 21.626 2.842 20.748 2.909 19.260 2.827
Electricity Demand 0.280 0.396 0.254 0.358 0.266 0.377 0.416 0.496 0.364 0.468 0.516 0.564 0.532 0.566 0.296 0.395
TaTS Benefits from Better Alignment (i.e., lower TT-Wasserstein). We compute the average forecasting improvement of TaTS compared to numerical-only modeling, as well as the ratio of TT-Wasserstein between original and shuffled Time-MMD datasets. The results, shown in Table 5, reveal that, within the same sampling frequency, lower TT-Wasserstein original-shuffled ratios tend to correlate with larger performance gains from TaTS (except for Climate). In other words, for paired time series and texts that have stronger CTR, i.e., more relevant, our TaTS could improve performance more. Thus, TT-Wasserstein can indicate the usefulness of paired texts and the potential effectiveness of TaTS when text quality and cross-modal alignment are uncertain.
Table 5: TaTS Improvements positively correlated with TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets. ”Original-Shuffled $T$ Ratio” means the ratio of TT-Wasserstein on the original dataset to TT-Wasserstein on the shuffled dataset. ”TaTS Improvement” means the average improvement of TaTS over uni-modal time series modeling.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
Original-Shuffled TT Ratio (%) 26.8 72.4 22.3 91.5 38.2 34.6 97.1 85.6 83.6
TaTS Improvement (%) 20.70 18.03 64.80 4.05 10.99 16.78 3.62 19.51 36.00
Comparison beyond Uni-modal Modeling and MM-TSFLib. Besides numerical-only time series models and the MM-TSFLib framework, we compare our TaTS with several other baselines on three dataset sources, including Time-MMD (Liu et al., 2024a), FNSPID (Dong et al., 2024), and FNF (Wang et al., 2024b). Although N-BEATS, N-HiTS and TCN are not specifically designed for time series with paired texts, we adapt them by feeding both time series and text embeddings to them. We also include several baselines that explicitly incorporate multimodal information.. For ChatTime, we concatenate all paired texts into a single long text and perform zero-shot inference. For GPT4MTS, we convert the datasets we used into their prescribed input format and directly apply their pipeline. The results averaged from multiple prediction lengths are shown in Table 4, with full results in Table 28. The results show that covariate-based and convolution-based models perform comparably or worse than iTransformer, and consistently underperform compared to our TaTS. Although ChatTime achieves competitive results even in a zero-shot setting, highlighting the value of integrating texts when high-quality texts are available, our TaTS outperforms it through supervised learning. We further compare TaTS with several multimodal methods from concurrent preprints that also evaluate on Time-MMD. Using their reported results (Table 12), we find that TaTS elevates standard time-series backbones into highly competitive models, rivaling recent approaches with substantially more complex designs.
5.2 Further Analysis
Hyperparameter Sensitivity. We perform hyperparameter studies to evaluate the impact of (i) the learning rate and (ii) $d_{\text{mapped}}$, the dimension to which high-dimensional text embeddings are projected by the MLP, as defined in Equation 5. The results are presented in Figure 4, subfigures (a) and (b), with full results available in Appendix E.5 and Appendix E.6. The findings indicate that TaTS maintains robust performance across different choices of the learning rate and the text projection dimension $d_{\text{mapped}}$.
Ablation with Different Text Encoders in TaTS. While GPT-2 was used as the primary text encoder in our main experiments to demonstrate the effectiveness of TaTS, we further evaluate the performance of TaTS with different text encoders, including BERT (Devlin et al., 2019), and LLaMA2 (Touvron et al., 2023). We utilize the official implementations available on Hugging Face and present the results in Figure 4, with full results provided in Appendix E.7. The results show that TaTS remains robust across different text encoders and consistently outperforms both the uni-modal and MM-TSFLib baselines. Notably, as the size of the language models used in TaTS increases from 110M (BERT) to 1.5B (GPT-2) and further to 7B (LLaMA2), we observe a slight improvement in performance. Investigating the relationship between the text encoder size and TaTS’s effectiveness remains an open direction for future research.
Table 6: TaTS Ablation on Alternative Modality Fusion Architectures.
Dataset Climate Security Traffic # Parameters
TaTS Settings MSE MAE MSE MAE MSE MAE Fuser Total
MLP projection (original) 0.992 0.791 109.865 4.968 0.178 0.224 74988 6396658
gated residual 0.998 0.794 110.254 4.965 0.177 0.227 9396 6331066
cross-attention 0.989 0.792 109.729 4.976 0.178 0.221 99340 6421010
Ablation with Other Design Choices to Combine Modalities. In our main experiments, TaTS transforms the paired texts into auxiliary variables through embedding, projection and concatenation. Our TaTS framework is readily extensible to other fusion design choices. We evaluate two alternative architectures that replace the MLP projection with: (a) a gated residual and (b) a cross-attention module between the time series and text embeddings. We report the average MSE and MAE across TaTS with iTransformer, PatchTST, and FiLM in Table 6. From the results, these alternative mechanisms achieve similar performance to the MLP-based design. One possible explanation is that linear projections are already highly competitive for time-series representation learning, especially when paired with strong backbone forecasters that account for the majority of the model parameters. We leave the exploration of more fine-grained multimodal fusion designs as an interesting direction for future work.
Table 7: TaTS Ablation on text-shuffled datasets.
Dataset Climate Security Traffic
Settings MSE MAE MSE MAE MSE MAE
iTransformer + TaTS + original data 1.028 0.804 112.05 5.151 0.187 0.217
iTransformer + TaTS + corrupted data 1.242 0.895 117.82 5.767 0.223 0.265
iTransformer + uni-modal + original data 1.183 0.871 115.94 5.660 0.213 0.238
Ablation on Corrupted Datasets. To validate that the performance gains stem from useful textual cues, we conduct corruption experiments where the textual information was randomly shuffled across timestamps. All other settings are unchanged, and the results are in Table 7. When textual alignment is destroyed, the performance drops significantly to matching or even being worse than the uni-modal baseline. This is expected, as randomly shuffled text acts as noise rather than a meaningful signal. Yet our TaTS stays relatively robust because it can learn to assign a small weight to textual information during optimization as a mitigation to noisy texts.
Table 8: TaTS Ablation on text-dropped datasets.
Dataset Climate Security Traffic
Settings MSE MAE MSE MAE MSE MAE
TaTS + original data 1.028 0.804 112.05 5.151 0.187 0.217
TaTS + 10% text randomly dropped 1.032 0.809 112.94 5.372 0.194 0.239
TaTS + 25% text randomly dropped 1.056 0.818 114.59 5.231 0.203 0.254
Ablation on Text-Randomly-Dropped Datasets. Similar to corruption ablations, we conduct experiments with randomly dropped texts. Dropped texts are filled with “no information available.” The results are in Table 8. While performance degrades slightly as more text is randomly dropped, TaTS remains effective, performing comparably to MM-TSF Lib even with 25% of the text missing. This experiment shows that whenever textual information is available, our TaTS could benefit from it to improve time series modeling.
Table 9: Dropping extremely noisy texts mitigates negative effects from them.
Dataset Climate Security Traffic
Settings MSE MAE MSE MAE MSE MAE
TaTS + corrupted data 1.242 0.895 117.82 5.767 0.223 0.265
TaTS + 60% corrupted data 1.218 0.879 117.43 5.730 0.215 0.260
TaTS + 20% corrupted data 1.201 0.878 116.68 5.713 0.216 0.255
uni-modal + original data 1.183 0.871 115.94 5.660 0.213 0.238
Mitigating Negative Effects of Extremely Noisy Texts. Our ablation on corrupted datasets shows that highly noisy text can cause TaTS to even slightly underperform numerical-only modeling. To address this, we evaluate a simple mitigation strategy: randomly dropping a portion of the noisy texts and replacing them with “no information available.” We test drop rates of 40% and 80%. As shown in Table 9, dropping noisy texts allows TaTS to recover performance close to the unimodal baseline, demonstrating the effectiveness of this strategy in handling unreliable text. Practitioners can refer to Appendix B.2 for how TaTS can adapt to scenarios with low-quality texts.
Figure 4: Further analysis of our TaTS framework. (a) Learning rate sensitivity: TaTS maintains stable performance across different learning rates (full results in Appendix E.5). (b) Text Projection Dimension sensitivity: TaTS remains robust across varying $d_{\text{mapped}}$ (full results in Appendix E.6). (c) Varying text encoder: TaTS consistently outperforms baselines across different text encoders (full results in Appendix E.7). (d) Efficiency: TaTS introduces only a minor parameter increase ($\sim 1\%$) but significantly improves the performance according to Table 2.
Computational Overhead vs. Performance Gain. We also evaluate the efficiency of our proposed TaTS by measuring the training time per epoch and the total number of model parameters. Figure 4 (d) presents the total number of parameters for the best-performing models in our forecasting experiments. TaTS introduces only a lightweight three-layer MLP to project high-dimensional text embeddings into a lower-dimensional space, adding a minimal number of parameters compared to the original time series models. As a result, the overall parameter count increases by only about 1%.
Due to the inclusion of augmented time series with auxiliary variables from paired texts, the training time per epoch increases slightly, as shown in Figure 5, with average performance of each framework indicated by cross markers. Full results for all datasets are provided in Appendix E.8. Notably, this marginal efficiency trade-off ($\sim 1\%$ in terms of number of learnable parameters and $\sim 8\%$ in terms of training time) leads to significant improvements ($\sim 14\%$) in forecasting performance.
Figure 5: While TaTS incurs a slight increase in training time due to augmented auxiliary variables, it significantly improves forecasting. Full results in Appendix E.8.
6 Related Work
Numerical-only Time Series Modeling. Recently, various deep learning models have been developed for time series analysis, which can be broadly categorized into three categories. (1) Patch-based models. PatchTST (Nie et al., 2023) segments time series into subseries-level patches to capture dependencies, while Crossformer (Zhang and Yan, 2023) employs a two-stage attention mechanism to model both cross-time and cross-variable dependencies efficiently. Autoformer (Wu et al., 2021) introduces decomposition blocks to separate seasonal and trend-cyclical components. (2) Global representation models. iTransformer (Liu et al., 2024c) utilizes attention over global series representations to capture multivariate correlations. Informer (Zhou et al., 2021) reduces self-attention complexity using ProbSparse self-attention for improved efficiency. Dlinear (Zeng et al., 2023) demonstrates that simple linear regression in the raw space can perform competitively on MTS tasks. (3) Frequency-aware models. FEDformer (Zhou et al., 2022b) represents series through randomly selected Fourier components, while FiLM (Zhou et al., 2022a) enhances representations with frequency-based layers to reduce noise and accelerate training. In this work, our proposed TaTS is compatible with all of the models listed above.
Time Series with other data sources. In the financial domain, several early works have explored integrating time series with textual data, albeit not in a timestamp-aligned manner, or often leveraging general machine learning models rather than time series-specific architectures. For example, StockNet (Xu and Cohen, 2018) uses a VAE-like model for chaotic stock-text data, while (Rodrigues et al., 2019) fuses time series with a single event-related document. BoEC (Farimani et al., 2021) applies a bag-of-economic-concepts approach, and Dandelion (Zhou et al., 2020) leverages multimodal attention for feature aggregation from multiple text sources. Some works also explored time series with vision information (Liu and Cai, 2012; Gerard et al., 2023; Lütjens et al., 2024) for spatio-temporal analysis. Recently, Time-MMD (Liu et al., 2024a) constructs a dataset of time series paired with parallel text, covering multiple domains, which we used in our main experiments.
Large Language Models on Time Series. The rapid advancement of Large Language Models (LLMs) has inspired a new line of research that transforms time series into natural language representations, enabling LLMs to perform downstream tasks (Zhang et al., 2024c). While these approaches demonstrate strong generalization capabilities due to the power of LLMs (Gruver et al., 2023; Cao et al., 2024; Jin et al., 2024; Xue and Salim, 2024), they also inherit limitations such as hallucination (Huang et al., 2023) and context length constraints (Wang et al., 2024a; Liu et al., 2024b). Notably, a recent work (Tan et al., 2024) suggests that replacing complex LLM architectures with basic attention layers does not necessarily degrade the performance. Another emerging line of work focuses on generating texts from time series, with applications in temporal reasoning and time-series question answering (Chang et al., 2025). MTBench (Chen et al., 2025) provides a comprehensive evaluation framework for assessing whether LLMs can jointly reason over structured numerical trends and unstructured textual narratives. TSAIA (Ye et al., 2025) benchmarks LLMs’ multi-step temporal reasoning capabilities. TimeXL (Jiang et al., 2025) further introduces collaborating LLM agents to generate interpretable natural-language explanations alongside time-series forecasts. One notable future direction is to enable latent communications (Zou et al., 2025a; b) between the time series model and language model to build general time series intelligence.
7 Conclusion
Real-world time series data often comes with textual descriptions, yet prior studies have largely overlooked this modality. We identify Chronological Textual Resonance, where text embeddings exhibit periodic patterns similar to their paired time series. To leverage this insight, we propose a plug-and-play framework that transforms text representations into auxiliary variables, seamlessly integrating them into existing time series models. Extensive experiments across various forecasting models and real-world datasets demonstrate the state-of-the-art performance of our approach.
Appendix
Roadmap. In this appendix, we provide a detailed overview of our methodology and experimental setup. Appendix A outlines the complete process of frequency analysis for both time series and paired texts. Appendix B discusses of preprints that are related to this work. Appendix C illustrates the TaTS algorithms for time series forecasting and imputation. Appendix D includes details on datasets, hyperparameters, evaluation metrics, and additional implementation specifics. Due to space constraints in the main text, Appendix E presents the full experimental results, including comprehensive forecasting and imputation outcomes, hyperparameter and ablation studies, efficiency evaluations, and visualizations. The table of contents is provided below for reference.
Appendix A Detailed Frequency Analysis Process of Time Series with Paired Texts
Here, we provide a detailed explanation of the frequency analysis process for both the time series and their paired texts. In Proposition A.1 and A.2, we simplify the proof by analyzing each periodic component individually as a single cosine. This is without loss of generality, as the operations involved are linear. Therefore, the conclusions naturally extend to signals composed of multiple periodic components through linear superposition.
Proposition A.1.
The computation of lag similarity preserves the original periodicities of the data.
Proof.
Let $\mathbf{S}=\{s_{t}\}_{t=1}^{T}$ represent the paired texts or data sequence, where each $s_{t}$ corresponds to a time step $t$. Define the lag similarity at lag $k$ as:
$\text{LagSim}(k)=\frac{1}{T-k}\sum_{t=1}^{T-k}\text{sim}(s_{t},s_{t+k}),$ (9)
where $\text{sim}(\cdot,\cdot)$ is a similarity measure (e.g., cosine similarity).
Now, consider the periodic component of the data sequence $\mathbf{S}$, which can be represented as:
$s_{t}=A\cos\left(\frac{2\pi t}{P}+\phi\right),$ (10)
where $A$ is the amplitude, $P$ is the period, and $\phi$ is the phase.
For two points separated by lag $k$, the similarity $\text{sim}(s_{t},s_{t+k})$ depends on the relative difference between their phases:
$s_{t+k}=A\cos\left(\frac{2\pi(t+k)}{P}+\phi\right)=A\cos\left(\frac{2\pi t}{P}+\frac{2\pi k}{P}+\phi\right).$ (11)
The lag similarity is then computed as:
$\text{LagSim}(k)=\frac{1}{T-k}\sum_{t=1}^{T-k}\text{sim}\left(A\cos\left(\frac{2\pi t}{P}+\phi\right),A\cos\left(\frac{2\pi t}{P}+\frac{2\pi k}{P}+\phi\right)\right).$ (12)
Since the cosine function is periodic with period $P$, the similarity $\text{sim}(s_{t},s_{t+k})$ also inherits this periodicity. Therefore, the overall lag similarity $\text{LagSim}(k)$ retains the periodicities of the original sequence $\mathbf{S}$.
Thus, the computation of lag similarity preserves the original periodicities of the data. ∎
Proposition A.2.
The stabilization of a data sequence using first-order differentiation preserves its original periodicities.
Proof.
Let $\mathbf{S}=\{s_{t}\}_{t=1}^{T}$ represent a data sequence, where $s_{t}$ is the value at time step $t$. The first-order differentiation of the sequence is defined as:
$\Delta s_{t}=s_{t+1}-s_{t},\quad t=1,2,\ldots,T-1.$ (13)
Suppose the sequence $\mathbf{S}$ exhibits periodic behavior with period $P$ and can be represented as:
$s_{t}=A\cos\left(\frac{2\pi t}{P}+\phi\right),$ (14)
where $A$ is the amplitude, $P$ is the period, and $\phi$ is the phase.
The first-order difference of $s_{t}$ is:
$\Delta s_{t}=s_{t+1}-s_{t}=A\cos\left(\frac{2\pi(t+1)}{P}+\phi\right)-A\cos\left(\frac{2\pi t}{P}+\phi\right).$ (15)
Using the trigonometric identity for the difference of cosines:
$\cos(x+y)-\cos(x)=-2\sin\left(\frac{y}{2}\right)\sin\left(x+\frac{y}{2}\right),$ (16)
we set $x=\frac{2\pi t}{P}+\phi$ and $y=\frac{2\pi}{P}$, giving:
$\Delta s_{t}=-2A\sin\left(\frac{\pi}{P}\right)\sin\left(\frac{2\pi t}{P}+\phi+\frac{\pi}{P}\right).$ (17)
The first term, $\sin\left(\frac{\pi}{P}\right)$, is a constant dependent on the period $P$. The second term, $\sin\left(\frac{2\pi t}{P}+\phi+\frac{\pi}{P}\right)$, retains the periodicity of $P$, as it is a sinusoidal function with the same frequency as the original sequence.
Thus, the first-order difference $\Delta s_{t}$ preserves the periodicity $P$ of the original data sequence. ∎
Figure 6: Illustration of the frequency analysis process for time series with paired texts in the Economy dataset. Step ①: Compute the average text similarity for each lag $k$. Step ②: Stabilize the time series using first-order differentiation, apply Fourier Transform, and perform Non-Maximum Suppression (NMS) to obtain the amplitude spectrum. Step ③: Visualize the lag similarity of paired texts. Step ④: Stabilize the paired texts, compute the Fourier Transform, and visualize the amplitude spectrum. Step ⑤: Overlay the top-$l$ (here ll=4) frequencies of paired texts onto the time series amplitude spectrum to highlight shared periodic patterns. All the data transformation operations in this process are periodicity-preserving according to Proposition A.1 and Proposition A.2.
The overall process of frequency analysis for the time series $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}\in\mathbb{R}^{T\times N}$ and paired texts $\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}$ in the Economy dataset $\mathcal{D}_{\text{Economy}}=\{\mathbf{X},\mathbf{S}\}$ is illustrated in Figure 6. In this process, starting from the original dataset (subfigure ①), the time series and paired texts are analyzed independently. For the time series $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ , we perform standard frequency analysis, stabilizing the data through first-order differentiation.
$\Delta\mathbf{X}_{t}=\mathbf{X}_{t+1}-\mathbf{X}_{t},\quad\text{for }t=1,2,\dots,T-1,$ (18)
where $\Delta\mathbf{X}_{t}$ represents the first-order difference of the time series. This step removes long-term trends and ensures that the data is stationary, allowing for a more accurate analysis of its frequency components.
Then, we compute the Fourier Transform (Nussbaumer, 1982; Sneddon, 1995) of $\Delta\mathbf{X}$ to analyze its frequency components. The Fourier Transform of $\Delta\mathbf{X}$ is defined as:
$\mathcal{F}_{\Delta\mathbf{X}}(f)=\sum_{t=1}^{T-1}\Delta\mathbf{X}_{t}e^{-i2\pi ft},$ (19)
where $f$ represents the frequency, $\Delta\mathbf{X}_{t}$ is the first-order difference of the time series at time $t$, and $i$ is the imaginary unit. The resulting $\mathcal{F}_{\Delta\mathbf{X}}(f)$ provides the amplitude and phase information of each frequency component present in the time series.
The magnitude spectrum, which represents the amplitude of each frequency component, is computed as:
$|\mathcal{F}_{\Delta\mathbf{X}}(f)|=\sqrt{\text{Re}(\mathcal{F}_{\Delta\mathbf{X}}(f))^{2}+\text{Im}(\mathcal{F}_{\Delta\mathbf{X}}(f))^{2}},$ (20)
where $\text{Re}(\mathcal{F}_{\Delta\mathbf{X}}(f))$ and $\text{Im}(\mathcal{F}_{\Delta\mathbf{X}}(f))$ are the real and imaginary parts of $\mathcal{F}_{\Delta\mathbf{X}}(f)$, respectively.
By analyzing $|\mathcal{F}_{\Delta\mathbf{X}}(f)|$, we identify the dominant frequencies in the time series, which reveal its periodic patterns. To further highlight the dominant frequencies, we apply Non-Maximum Suppression (NMS) to the magnitude spectrum $|\mathcal{F}_{\Delta\mathbf{X}}(f)|$. NMS ensures that only the most prominent frequencies are retained while suppressing nearby less significant frequencies. The NMS operation is defined as follows:
$\mathcal{N}(f)=\begin{cases}|\mathcal{F}_{\Delta\mathbf{X}}(f)|,&\text{if }|\mathcal{F}_{\Delta\mathbf{X}}(f)|>|\mathcal{F}_{\Delta\mathbf{X}}(f^{\prime})|\,\forall f^{\prime}\in\mathcal{N}(f),\\ 0,&\text{otherwise,}\end{cases}$ (21)
where $\mathcal{N}(f)$ represents a local neighborhood around the frequency $f$. The operation compares the magnitude of $|\mathcal{F}_{\Delta\mathbf{X}}(f)|$ with those of neighboring frequencies and retains only the largest value within the neighborhood. Frequencies that do not satisfy the condition are set to zero.
After applying NMS, the remaining frequencies represent the dominant periodic components of the time series, making it easier to identify significant periodic patterns. This process eliminates noise and reduces the influence of minor frequency components, enhancing the interpretability of the spectrum. The final visualization of the amplitude spectrum of the time series is shown in Figure 6, subfigure ②.
For the paired texts $\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}$, we first embed each $s_{t}$ using the text encoder $\mathcal{H}_{\text{text}}$:
$\mathbf{e}_{t}=\mathcal{H}_{\text{text}}(s_{t};\theta_{\text{text}})\in\mathbb{R}^{d_{\text{text}}},$ (22)
where $\theta_{\text{text}}$ represents the parameters of the text encoder, and $\mathbf{e}_{t}$ is the resulting text embedding at timestamp $t$.
Since the text embeddings are typically close in the embedding space, leading to similar cosine similarity values, we normalize the embeddings by centering them around their mean to improve numerical stability and enhance sensitivity to differences. Specifically, we compute the mean embedding:
$\mathbf{e}_{\text{mean}}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{e}_{t},$ (23)
and shift all embeddings by subtracting the mean:
$\mathbf{e}_{t}^{\prime}=\mathbf{e}_{t}-\mathbf{e}_{\text{mean}},$ (24)
where $\mathbf{e}_{t}^{\prime}$ represents the centered (shifted) embeddings.
Next, we compute the average text similarity for each lag $k\in\{1,T-1\}$ as:
$\text{Sim}(k)=\frac{1}{T-k}\sum_{t=1}^{T-k}\text{sim}(\mathbf{e}_{t}^{\prime},\mathbf{e}_{t+k}^{\prime}),$ (25)
where $\text{sim}(\mathbf{e}_{t}^{\prime},\mathbf{e}_{t+k}^{\prime})$ denotes the similarity measure (e.g., cosine similarity) between the centered embeddings at time $t$ and $t+k$, defined as:
$\text{sim}(\mathbf{e}_{t}^{\prime},\mathbf{e}_{t+k}^{\prime})=\frac{\mathbf{e}_{t}^{\prime}\cdot\mathbf{e}_{t+k}^{\prime}}{\|\mathbf{e}_{t}^{\prime}\|\|\mathbf{e}_{t+k}^{\prime}\|}.$ (26)
We visualize the lag similarity of paired texts, $\text{Sim}(k)$, in Figure 6, subfigure ③. Subsequently, we stabilize the data by applying first-order differentiation and perform a Fourier Transform, following a similar process as previously described for the time series frequency analysis. The final visualization of the amplitude spectrum of the paired texts is presented in Figure 6, subfigure ④.
Then, we compute the frequencies with the top-$l$ amplitudes from the lag similarity $\text{Sim}(k)$. We apply the Fourier Transform to $\text{Sim}(k)$:
$\mathcal{F}_{\text{text}}(f)=\sum_{t=1}^{T-1}\text{Sim}(k)\,e^{-i2\pi ft},$ (27)
where $f$ is the frequency, and $\mathcal{F}_{\text{text}}(f)$ represents the complex Fourier coefficients corresponding to each frequency $f$.
Next, we compute the amplitude spectrum as:
$|\mathcal{F}_{\text{text}}(f)|=\sqrt{\text{Re}(\mathcal{F}_{\text{text}}(f))^{2}+\text{Im}(\mathcal{F}_{\text{text}}(f))^{2}},$ (28)
where $\text{Re}(\mathcal{F}_{\text{text}}(f))$ and $\text{Im}(\mathcal{F}_{\text{text}}(f))$ are the real and imaginary parts of $\mathcal{F}_{\text{text}}(f)$, respectively.
We then identify the top-$l$ dominant frequencies by selecting the $l$ frequencies corresponding to the largest amplitudes:
$\mathcal{F}_{\text{top}}=\{f_{i}\mid|\mathcal{F}_{\text{text}}(f_{i})|\text{ is among the top-}l\text{ largest amplitudes}\}.$ (29)
These top-$l$ frequencies represent the most significant periodic components of the paired texts, revealing their dominant temporal patterns. we overlay the top-$l$ (in the Economy dataset ll=4) frequencies of the paired texts onto the amplitude spectrum of the time series, as illustrated in Figure 6, subfigure ⑤.
We also visualize the frequency analysis process in the Social Good dataset and Traffic dataset respectively in Figure 7 and Figure 8.
Figure 7: Illustration of the frequency analysis process for time series with paired texts in the Social Good dataset. In Step ⑤, we overlay the top-$9$ frequencies of paired texts onto the time series amplitude spectrum.
Figure 8: Illustration of the frequency analysis process for time series with paired texts in the Traffic dataset. In Step ⑤, we overlay the top-$7$ frequencies of paired texts onto the time series amplitude spectrum.
A.1 Estimation Stability of TT-Wasserstein
As TT-Wasserstein is an empirically estimated statistical metric, we provide a detailed sensitivity and robustness analysis in this appendix. Our default configuration uses a rectangular window, $\ell_{1}$-normalization of spectral amplitudes, and a frequency grid determined by the full length of the time series (native FFT resolution). To assess the stability of TT-Wasserstein, we perform ablations that vary each of these components: windowing schemes, normalization strategies, and frequency resolutions, while keeping all other settings fixed. The results are summarized in Table 10, where each row reflects the effect of a single deviation from the default configuration. From the results, our TT-Wasserstein is stable across various configuration choices. We also implemented block bootstrap with a bootstrap sample size of 10 for time series and texts to compute the confidence intervals for the default settings in Table 11.
Table 10: TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets. Our TT-Wasserstein is generally stable on different frequency computation configurations.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
Default 0.026 0.025 0.022 0.049 0.027 0.035 0.307 0.233 0.302
Hann Window 0.031 0.029 0.016 0.041 0.040 0.057 0.304 0.243 0.306
Hamming Window 0.054 0.025 0.016 0.041 0.037 0.046 0.305 0.242 0.305
Blackman Window 0.061 0.029 0.020 0.036 0.043 0.055 0.297 0.245 0.298
Minmax Normalization 0.032 0.035 0.025 0.043 0.031 0.049 0.345 0.291 0.332
Log Normalization 0.016 0.060 0.046 0.049 0.050 0.048 0.267 0.170 0.286
Zero-padded frequency resolution ($2\times$) 0.046 0.058 0.052 0.051 0.028 0.050 0.280 0.187 0.292
Down-sampled frequency resolution ($\frac{1}{2}\times$) 0.011 0.026 0.039 0.050 0.051 0.025 0.259 0.197 0.298
Table 11: Standard deviation of default TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets by block bootstrap.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
Default $\pm$0.06 $\pm$0.04 $\pm$0.03 $\pm$0.03 $\pm$0.04 $\pm$0.06 $\pm$0.21 $\pm$0.16 $\pm$0.12
Appendix B Further Discussion
B.1 Concurrent Works
We discuss concurrent preprints available online that relate to this research, providing a more comprehensive understanding of existing works and ongoing efforts to integrate texts into time series modeling.
(Williams et al., 2024) introduces a benchmark (CiK) for time-series forecasting that integrates both numerical data and textual context. While both (Williams et al., 2024) and our work emphasize the integration of textual context into time series forecasting, the textual contexts in (Williams et al., 2024) are descriptions regarding the time series itself, but in our work, we do not have such a constraint.
(Kim et al., 2024) develops a hybrid forecaster that jointly predicts both time series and textual data by projecting time series to the language space and fine-tuning the pre-trained LLM. Their assumption is that though LLMs are originally for language but not for time series, they can be fine-tuned to adapt to time series.
(Zhang et al., 2024a) introduces a dual-adapter model for time series and textual data, but their data format is similar to (Wang et al., 2025), where the whole time series is paired with one text.
(Xie et al., 2024) proposes a novel multimodal LLM framework (TS-MLLM) that treats time series as a modality akin to images. But they are using synthetic generation and focus on NLP tasks rather than time series forecasting.
Comparison with Concurrent Multimodal Time Series Forecasting Methods. Several concurrent works also explore incorporating multimodal information to enhance time series forecasting and report results on the Time-MMD benchmark (Liu et al., 2024a). TFHTS (Zhou et al., 2025) adopts a dual-tower architecture to fuse time series and textual signals, followed by large language models for forecasting. MCD-TSF (Su et al., 2025b) proposes a multimodal-conditioned diffusion model that adaptively aligns textual context with temporal dynamics. TeR-TSF (Su et al., 2025a) introduces text reinforcement, generating augmented textual descriptions to improve downstream prediction.
For a fair comparison, we follow their evaluation protocol and exclude datasets that exhibit substantial performance gaps, suggesting that their preprocessing pipelines may differ from ours. we report averaged performance over prediction lengths 6, 12, and 18 for monthly-frequency datasets (Agriculture, Climate, Social Good); 12, 24, and 36 for weekly-frequency datasets (Energy, Health); and 48, 96, and 192 for the daily-frequency Environment dataset. Note that, unlike our main experiments, we use a historical window of 36 steps for monthly, 96 for weekly, and 192 for daily datasets. The results are summarized in Table 12. From the results, TaTS achieves the best or second-best performance on 9 out of the 12 metrics. To provide an overall comparison, we compute an average ranking as follows: for each dataset, we assign each method a rank from 1 to 4 for both MSE and MAE (with lower values indicating better performance), average the two ranks to obtain a per-dataset score, and then average these scores across all datasets. Under this ranking scheme, TaTS obtains the smallest average ranking, indicating the best overall performance among concurrent multimodal time series forecasting methods.
Table 12: TaTS compared with several concurrent works that have been evaluated on Time-MMD. We use the reported performance in the concurrent works for comparison. Best results are bolded and second-best results are underlined. Full results in Table 28.
Methods TaTS (Ours) TFHTS MCD-TSF TeR-TSF
+ iTransformer (2025) (2025b) (2025a)
Datasets
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Time-MMD: Multimodal Time Series (2024a) Agriculture 0.201 0.321 0.571 0.564 0.222 0.322 0.338 0.402
Climate 1.227 0.890 1.782 0.948 1.583 0.971 1.348 0.874
Energy 0.214 0.321 0.290 0.403 0.153 0.293 0.202 0.324
Environment 0.267 0.374 0.262 0.368 0.275 0.379 0.251 0.346
Health 1.372 0.763 1.514 0.809 1.496 0.811 1.384 0.756
Social Good 1.146 0.558 1.352 0.677 1.035 0.569 1.199 0.587
Average Ranking 1.83 3.5 2.58 2.08
B.2 Limitations
This work focuses on revealing and quantifying Chronological Textual Resonance (CTR) and designing the TaTS framework to leverage it. However, several limitations remain. First, we do not thoroughly investigate the data construction processes that may induce CTR, such as biases introduced during text collection or the choice of contextual information. Understanding how these factors affect CTR would provide deeper insights into the robustness and generalizability of our approach. Second, we do not analyze how text embeddings are utilized at the neural level within the TaTS framework. A more detailed study on how the model learns temporal patterns from paired texts could inform improvements in architecture design. Third, the effectiveness of TaTS is influenced by the quality and relevance of paired texts. While we study cases where texts are randomly shuffled, real-world texts could have different patterns, and the model’s performance may degrade. Future work could explore methods to assess and enhance text quality or develop more robust models to handle noisy input. Lastly, while TaTS demonstrates strong empirical performance across benchmark datasets, its generalization to other types of multimodal time series or domains with fundamentally different patterns remains untested, for example, time series with paired images or audio. We leave these investigations and improvements for future work.
Regarding our third limitation on text quality, we briefly clarify how TaTS may adapt under imperfect textual inputs. When texts are missing at some timestamps, TaTS can simply use a placeholder token (e.g., “no information available”), as validated in our text-random-drop ablations. When temporal alignment is uncertain, small lead–lag shifts mainly affect phase rather than frequency, so TaTS can still leverage the preserved periodic structure. For larger misalignments, standard preprocessing alignment methods may be applied, with TT-Wasserstein serving as a diagnostic to guide how strongly to rely on the text modality.
B.3 Ethical Impact Statement
Our work is solely focused on the technical challenge of multimodal time series and does not involve any elements that could pose ethical risks.
Appendix C Algorithms
C.1 Pseudo Code
1
Input: Time series with concurrent texts embeddings $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\}$ in the input training dataset; prediction length $H$.
Output: TaTS model parameters $\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}$.
2
3Prepare training samples of sequence length $L$ and prediction length $H$ as:
$\{\mathbf{X}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{E}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{Y}^{(i)}=\mathbf{X}_{l_{i}+L+1:l_{i}+L+H}\}_{i=1}^{n}$
Initialize time series model $\mathcal{F}(\cdot;\theta_{\text{forecast}})$; projector $\text{MLP}(\cdot;\theta_{\text{MLP}})$ for dimensionality reduction.
4while not converged do
5 for each training sample $\{\mathbf{X}^{(i)},\mathbf{E}^{(i)},\mathbf{Y}^{(i)}\}$ do
6      Map $\mathbf{E}^{(i)}$ to $\mathbf{Z}^{(i)}$: $\mathbf{Z}^{(i)}[j]=\text{MLP}(\mathbf{E}^{(i)}[j];\theta_{\text{MLP}})$
7     Compute $\mathbf{U}=[\mathbf{X}^{(i)};{(\mathbf{Z}^{(i)})}^{\intercal}]_{\text{dim=1}}$ as shown in Eq. (6).
8     Forecast: $\widehat{\mathbf{X}}^{(i)}=\mathcal{F}(\mathbf{U};\theta_{\text{forecast}})[:N]$
9     Optimize: $\operatorname*{arg\,min}_{\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}}\mathcal{L}_{\text{forecast}}(\mathbf{X}^{(i)},\widehat{\mathbf{X}}^{(i)})$ as shown in Eq. (8)
10   end for
11
12 end while
13
return TaTS model parameters $\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}$
Algorithm 1 Texts as Time Series for Forecasting Task
1
Input: Time series with concurrent text embeddings $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\}$ in the input training dataset; binary mask $\mathbf{M}\in\{0,1\}^{T\times N}$.
Output: TaTS model parameters $\Theta=\{\theta_{\text{impute}},\theta_{\text{MLP}}\}$.
2
3Prepare training samples with observed entries:
$\{\mathbf{X}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{E}^{(i)}=\mathbf{E}_{l_{i}+1:l_{i}+L},\mathbf{M}^{(i)}=\mathbf{M}_{l_{i}+1:l_{i}+L}\}_{i=1}^{n}$
Initialize time series imputation model $\mathcal{G}(\cdot;\theta_{\text{impute}})$; projector $\text{MLP}(\cdot;\theta_{\text{MLP}})$ for dimensionality reduction.
4while not converged do
5 for each training sample $\{\mathbf{X}^{(i)},\mathbf{E}^{(i)},\mathbf{M}^{(i)}\}$ do
6      Map $\mathbf{E}^{(i)}$ to $\mathbf{Z}^{(i)}$: $\mathbf{Z}^{(i)}[j]=\text{MLP}(\mathbf{E}^{(i)}[j];\theta_{\text{MLP}})$
7     Construct the augmented input: $\mathbf{U}=[(\mathbf{X}^{(i)}\odot\mathbf{M}^{(i)});{(\mathbf{Z}^{(i)})}^{\intercal}]_{\text{dim=1}}$
8     Impute missing values: $\widehat{\mathbf{X}}^{\text{Imputed}(i)}=\mathcal{G}(\mathbf{U};\theta_{\text{impute}})$
9     Optimize: $\operatorname*{arg\,min}_{\Theta=\{\theta_{\text{impute}},\theta_{\text{MLP}}\}}\mathcal{L}_{\text{impute}}(\mathbf{X}^{(i)},\widehat{\mathbf{X}}^{\text{Imputed}(i)})$
10   end for
11
12 end while
13
return TaTS model parameters $\Theta=\{\theta_{\text{impute}},\theta_{\text{MLP}}\}$
Algorithm 2 Texts as Time Series for Imputation Task
C.2 Incremental Complexity Introduced by the Text Modality
As analyzed in Section 5.2, incorporating textual information introduces some additional computational overhead. This overhead arises from two components: (1) the cost of encoding the text tokens, and (2) the cost of training the forecasting model on the augmented input. We have empirically evaluated the latter in Section 5.2; here, we provide a theoretical analysis of the former.
Let the language encoder process each token in time $O(t_{\text{token}})$, and let each timestamp have an associated text description with an average length of $l$ tokens. If the time series contains $T$ timestamps, then the total time complexity for embedding all paired texts is:
$O(t_{\text{token}}\cdot l\cdot T).$
This follows because each of the $T$ timestamps requires encoding, on average, $l$ tokens, and each token requires $O(t_{\text{token}})$ time to process. In practice, from our experiments reported in Table 13, this cost is modest for lightweight text encoders (e.g., small Transformers or MLP-based tokenizers), and the embedding step is performed only once per dataset rather than per training iteration. As a result, the added modality introduces only a manageable overhead relative to the overall cost of training modern time-series forecasting models.
Table 13: Wall-clock costs (seconds) to embed texts for Time-MMD (Liu et al., 2024a) datasets.
Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
BERT 0.12 0.21 0.23 0.23 1.7 0.21 0.26 0.46 0.29
GPT-2 0.18 0.25 0.31 0.28 2.5 0.30 0.28 0.55 0.35
LLaMA-2-7B 0.28 0.29 0.37 0.35 10.5 0.40 0.32 0.62 0.42
Appendix D Experiment Details
D.1 Dataset Statistics and Details
Table 14, Table 15 and Table 16 provide summaries of the statistics for the publicly available real-world datasets. Additionally, we visualize the numerical data for each Time-MMD multimodal time series dataset in Figure 9. For further details, please refer to the original work that introduced these datasets and benchmarks (Liu et al., 2024a).
(a) Agriculture
(b) Climate
(c) Economy
(d) Energy
(e) Environment
(f) Health
(g) Security
(h) Social Good
(i) Traffic
Figure 9: Visualization of the numerical data in each multimodal time series dataset.
Table 14: Overview of the numerical data in the Time-MMD datasets (Liu et al., 2024a). “Prediction Length” refers to the number of future time points to be forecasted, with each dataset including four distinct prediction settings. refers to the number of variables (or variates) in each dataset.
Dataset Name/Domain Prediction Length Dimension Frequency Number of Samples Timespan
Agriculture {6, 8, 10, 12} 1 Monthly $496$ 1983 - Present
Climate {6, 8, 10, 12} 5 Monthly $496$ 1983 - Present
Economy {6, 8, 10, 12} 3 Monthly $423$ 1989 - Present
Energy {12, 24, 36, 48} 9 Weekly $1479$ 1996 - Present
Environment {48, 96, 192, 336} 4 Daily $11102$ 1982 - 2023
Health {12, 24, 36, 48} 11 Weekly $1389$ 1997 - Present
Security {6, 8, 10, 12} 1 Monthly $297$ 1999 - Present
Social Good {6, 8, 10, 12} 1 Monthly $900$ 1950 - Present
Traffic {6, 8, 10, 12} 1 Monthly $531$ 1980 - Present
Table 15: Overview of the numerical data in the FNSPID (Financial News and Stock Price Integration Dataset) datasets (Dong et al., 2024). “Prediction Length” refers to the number of future time points to be forecasted, with each dataset including four distinct prediction settings. refers to the number of variables (or variates) in each dataset.
Company/Stock Name Prediction Length Dimension Frequency Number of Samples Timespan
Delta Airlines (DAL) {6, 8, 10, 12} 1 Bi-daily $1358$ 2009 - 2020
IBM (IBM) {6, 8, 10, 12} 1 Bi-daily $493$ 2016 - 2020
JPMorgan Chase (JPM) {6, 8, 10, 12} 1 Bi-daily $565$ 2018 - 2020
NVIDIA (NVDA) {6, 8, 10, 12} 1 Bi-daily $1203$ 2011 - 2020
Pfizer (PFE) {6, 8, 10, 12} 1 Bi-daily $812$ 2016 - 2020
Tesla (TSLA) {6, 8, 10, 12} 1 Bi-daily $294$ 2019 - 2020
Table 16: Overview of the numerical data in the FNF (From News to Forecast) datasets (Wang et al., 2024b). “Prediction Length” refers to the number of future time points to be forecasted, with each dataset including four distinct prediction settings. refers to the number of variables (or variates) in each dataset.
Dataset Name/Domain Prediction Length Dimension Frequency Number of Samples Timespan
Bitcoin Price {6, 8, 10, 12} 1 Daily $1237$ 2018 - 2021
Web Traffic {6, 8, 10, 12} 1 Daily $728$ 2015 - 2016
Electricity Demand {6, 8, 10, 12} 1 Daily $1097$ 2019 - 2021
D.2 Hyperparameters
We use Adam optimizer (Kingma and Ba, 2015) when training the neural networks. The default choices of hyperparameters in our code are provided in Table 17. For LLM-based text encoders, we initialize them using the default configurations provided by Hugging Face [^1] https://huggingface.co/. Consistent with existing works (Wu et al., 2023a; Liu et al., 2024c), we apply instance normalization to standardize the time series data within each dataset.
Table 17: Default hyperparameters for the TaTS framework
Hyperparameter Description Value or Choices
batch_size The batch size for training 32
criterion The criterion for calculating loss Mean Square Error (MSE)
learning_rate The learning rate for the optimizer {0.0001, 0.00005, 0.00001}
seq_len Input sequence length 24
label_len Start token length for prediction 12
prior_weight Weight for prior combination {0, 0.1, 0.2, 0.3, 0.5}
train_epochs Number of training epochs 50
patience Early stopping patience 20
text_emb Dimension of text embeddings {6, 12, 24}
learning_rate2 Learning rate for MLP layers {0.005, 0.01, 0.02, 0.05}
pool_type Pooling type for embeddings “avg”
init_method Initialization method for combined weights “normal”
dropout dropout 0.1
use_norm whether to use normalize True
D.3 Metrics
Throughout this paper, we use the following metrics to evaluate performance:
MSE (Mean Squared Error): Measures the average squared difference between the predicted and actual values. It penalizes larger errors more heavily, making it sensitive to outliers.
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (30)
where $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted values, respectively, and $n$ is the number of data points.
MAE (Mean Absolute Error): Represents the average absolute difference between the predicted and actual values, providing a more interpretable measure of average error magnitude.
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|.$ (31)
RMSE (Root Mean Squared Error): The square root of MSE, which provides an error measure in the same units as the original data. It is more sensitive to large deviations than MAE.
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2}}.$ (32)
MAPE (Mean Absolute Percentage Error): Expresses errors as a percentage of the actual values, offering a scale-independent metric that facilitates comparisons across datasets.
$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100.$ (33)
MSPE (Mean Squared Percentage Error): Similar to MAPE but squares the percentage error, penalizing larger percentage deviations more heavily.
$\text{MSPE}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right)^{2}.$ (34)
These metrics collectively provide a comprehensive evaluation of model performance, capturing both absolute and relative errors as well as their sensitivity to outliers. For all metrics, lower values indicate better performance.
D.4 Implementation Details
D.4.1 Code and Reproducibility
The code for the experiments is included in the supplementary material, accompanied by a comprehensive README file. We provide detailed commands, scripts, and instructions to facilitate running the code. Additionally, the datasets used in the experiments are provided in the supplementary material as CSV files.
D.4.2 Hardware and Environment
We conducted all experiments on an Ubuntu 22.04 machine equipped with an Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz, 1.5TB of RAM, and a 32GB NVIDIA V100 GPU. The CUDA version used was 12.4. All algorithms were implemented in Python (version 3.11.11). To run our code, users must install several commonly used libraries, including pandas, scikit-learn, patool, tqdm, sktime, matplotlib, transformers, and others. Detailed installation instructions can be found in the README file within the code directory. We have optimized our code to ensure efficiency. Our tests confirmed that the CPU memory usage remains below 16 GB, while the GPU memory usage is under 20 GB. Additionally, the execution time for a single experiment is less than 10 minutes on our machine.
D.4.3 Data Splitting and Leakage Prevention
To ensure forecasting without access to contemporaneous or future information, we adopt a chronological data-splitting protocol for all datasets. Timestamps are divided into $80\%$ training, $10\%$ validation, and $10\%$ test windows without shuffling. For a forecast at timestamp $T$, the model receives only time-series values and textual inputs with timestamps $\leq T-1$, while all inputs with timestamps $\geq T$ are masked during prediction.
Timestamp Alignment of Texts.
All datasets provide pre-aligned text–time-series pairs by associating each text sample with the timestamp at which it originally became publicly available (e.g., publication or release time). We directly adopt these timestamps, ensuring that the textual inputs available to the model at time T1T-1 correctly reflect real-time accessibility.
Auditing for Retrospective Leakage.
Although we do not apply automated filtering procedures, we conduct human audits on random subsets of samples from each dataset to detect retrospective or outcome-summarizing leakage (e.g., texts describing events that occur after the associated timestamp). Across all datasets, we did not observe such leakage. Combined with causal masking, this protocol prevents the models from accessing contemporaneous or future information from either modality.
Appendix E Full Experiment Results
E.1 Why should we leverage CTR?
Table 18: Performance on time series forecasting task. Leveraging periodicity with a single-dimension feature can significantly reduce the prediction error.
Method Economy Social Good Traffic
MSE($\downarrow$) MAE($\downarrow$) MSE($\downarrow$) MAE($\downarrow$) MSE($\downarrow$) MAE($\downarrow$)
Uniformly Random $(+)$ 5.673 2.356 2.059 1.230 1.207 0.995
Uniformly Random $(\pm)$ 11.535 2.879 8.860 2.404 3.794 1.618
Normally Random 9.284 2.878 2.926 1.374 3.163 1.511
Exponentially Random 4.521 1.960 3.724 1.564 1.141 0.911
Using 1D Text only 1.995 1.404 1.315 0.853 0.714 0.797
Parallel text provides complementary information and expert knowledge that can significantly enhance the understanding of time series data. To demonstrate the benefits of utilizing periodicity in the text modality, we present an illustrative example in the univariate forecasting task, where the goal is to use $\mathbf{X}_{1:T}=\{\mathbf{x}_{1}\}\in\mathbb{R}^{T\times 1}$ to predict the next $H$ values, $\widehat{\mathbf{X}}_{T+1:T+H}$. We first concatenate the text embeddings to form $\mathbf{E}=[e_{1};e_{2};\ldots;e_{T}]_{\text{dim=1}}\in\mathbb{R}^{d_{\text{text}}\times T}$, then replace $\mathbf{x}_{1}$ with only the first dimension of the text embeddings to leverage very partial text periodicity, $\mathbf{x}^{\prime}_{1}=(\mathbf{E}[1,:])^{\intercal}\in\mathbb{R}^{T\times 1}$. In other words, we rely solely on a single evolving dimension of the paired text features to forecast future time series values.
As shown in Table 18, leveraging just the periodicity of a single text feature significantly outperforms random time series forecasting. The random baselines mean that the forecasts are purely random according to several different distributions. These random baselines use the training data as well, for example, normally random computes the mean and standard deviation of a normal distribution to forecast the time series. These results highlight that even partial periodicity from the text modality contributes valuable insights for improving forecasting accuracy.
E.2 Full Forecasting Performance Comparison Visualization
To provide a comprehensive comparison of different frameworks for modeling time series with paired texts, we visualize the forecasting performance using radar plots in Figure E.2. Each subfigure corresponds to a dataset, with each axis representing a different time series model. The axes are inverted, where values closer to the center indicate worse performance, and larger areas signify better results. The results demonstrate that TaTS consistently outperforms both baselines across all datasets while maintaining compatibility with various time series models.
(a) Agriculture
(b) Climate
(c) Economy
(d) Energy
(e) Environment
(f) Health
(g) Security
(h) Social Good
(i) Traffic
Figure 10: Comparison of different frameworks for modeling time series with paired texts. Our TaTS achieves the best performance across all datasets and is compatible with various existing time series models.
E.3 Full Forecasting Results
Due to space limitations, we provide the full results of the time series forecasting task on paired time series and text in the appendix. We conduct extensive experiments across 9 datasets using 9 existing time series models, evaluating various prediction lengths as detailed in Table 14. The complete results are presented from Table 19 to Table 27. Overall, TaTS consistently achieves the best performance across all datasets, time series models, and prediction lengths. The averaged results across all prediction lengths are summarized in the main text (Table 2). For better readability, we also visualize the performance of different frameworks using radar plots, as detailed in Appendix E.2 and Figure 10.
Table 19: Full forecasting results for the Agriculture, Climate, and Economy datasets using iTransformer, PatchTST, and Crossformer as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models iTransformer PatchTST Crossformer
(2024c) (2023) (2023)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Agriculture Uni-modal 6 0.077 0.200 0.274 0.090 0.014 0.074 0.197 0.269 0.090 0.014 0.222 0.331 0.412 0.136 0.031
8 0.104 0.228 0.315 0.100 0.017 0.104 0.234 0.317 0.104 0.018 0.304 0.406 0.496 0.168 0.044
10 0.142 0.273 0.372 0.119 0.024 0.136 0.262 0.357 0.112 0.021 0.357 0.435 0.530 0.176 0.049
12 0.167 0.301 0.400 0.128 0.026 0.168 0.294 0.396 0.124 0.025 0.409 0.451 0.582 0.181 0.054
Avg 0.122 0.251 0.340 0.109 0.020 0.120 0.247 0.335 0.107 0.020 0.323 0.406 0.505 0.165 0.044
$M$-TSFLib 6 0.070 0.189 0.261 0.085 0.013 0.071 0.183 0.261 0.082 0.012 0.146 0.259 0.331 0.106 0.020
8 0.091 0.212 0.296 0.093 0.016 0.093 0.212 0.296 0.093 0.015 0.171 0.278 0.354 0.112 0.022
10 0.130 0.248 0.346 0.105 0.019 0.126 0.251 0.344 0.108 0.020 0.252 0.344 0.428 0.136 0.032
12 0.158 0.272 0.377 0.112 0.022 0.168 0.288 0.391 0.119 0.024 0.304 0.372 0.471 0.144 0.036
Avg 0.112 0.230 0.320 0.099 0.018 0.114 0.233 0.323 0.100 0.018 0.218 0.313 0.396 0.124 0.027
TaTS (ours) 6 0.067 0.184 0.256 0.083 0.012 0.066 0.171 0.247 0.076 0.011 0.148 0.264 0.332 0.108 0.020
8 0.094 0.210 0.297 0.091 0.015 0.096 0.217 0.300 0.094 0.016 0.197 0.298 0.374 0.119 0.026
10 0.122 0.251 0.341 0.109 0.020 0.126 0.260 0.349 0.113 0.021 0.216 0.315 0.397 0.124 0.028
12 0.153 0.271 0.374 0.112 0.022 0.166 0.292 0.394 0.123 0.025 0.289 0.371 0.466 0.145 0.036
Avg 0.109 0.229 0.317 0.099 0.017 0.114 0.235 0.323 0.101 0.018 0.212 0.312 0.392 0.124 0.027
Climate Uni-modal 6 1.127 0.843 1.052 2.549 50.662 1.259 0.915 1.122 3.168 105.890 1.159 0.852 1.076 3.036 148.840
8 1.191 0.876 1.088 3.208 134.908 1.208 0.878 1.099 2.778 64.950 1.104 0.829 1.051 2.600 123.988
10 1.215 0.885 1.100 3.169 123.266 1.218 0.894 1.103 2.746 55.831 1.127 0.829 1.057 3.246 193.108
12 1.199 0.879 1.091 2.752 65.916 1.197 0.891 1.092 3.177 115.617 1.105 0.837 1.047 3.322 194.844
Avg 1.183 0.871 1.083 2.920 93.688 1.220 0.895 1.104 2.967 85.572 1.124 0.837 1.058 3.051 165.195
$M$-TSFLib 6 1.031 0.787 1.013 2.298 38.389 1.003 0.800 1.001 2.652 57.709 0.995 0.758 0.996 2.393 51.345
8 1.039 0.809 1.017 2.598 49.903 1.012 0.790 1.005 2.387 42.258 1.016 0.773 1.007 2.824 107.865
10 1.049 0.817 1.020 2.547 55.014 1.036 0.813 1.016 2.293 28.894 0.999 0.779 0.997 2.740 95.790
12 1.057 0.828 1.025 2.894 87.976 1.071 0.822 1.032 2.965 95.248 0.997 0.777 0.994 2.240 43.968
Avg 1.044 0.810 1.019 2.584 57.821 1.030 0.806 1.014 2.574 56.027 1.002 0.772 0.998 2.549 74.742
TaTS (ours) 6 1.020 0.797 1.007 2.563 56.209 0.976 0.782 0.987 2.318 33.843 0.924 0.747 0.961 1.895 20.405
8 1.025 0.797 1.011 2.391 38.756 0.995 0.803 0.997 2.574 57.639 0.923 0.757 0.961 2.318 51.592
10 1.033 0.808 1.014 2.647 77.697 1.022 0.796 1.007 2.657 76.287 0.963 0.764 0.979 2.439 59.881
12 1.033 0.812 1.013 2.607 58.107 1.022 0.810 1.009 2.436 42.471 0.943 0.754 0.967 2.220 40.186
Avg 1.028 0.804 1.011 2.552 57.692 1.004 0.798 1.000 2.496 52.560 0.938 0.755 0.967 2.218 43.016
Economy Uni-modal 6 0.015 0.099 0.124 0.035 0.002 0.017 0.104 0.129 0.036 0.002 0.659 0.749 0.806 0.257 0.076
8 0.014 0.098 0.120 0.034 0.002 0.016 0.104 0.128 0.036 0.002 0.661 0.767 0.808 0.263 0.076
10 0.014 0.094 0.119 0.033 0.002 0.017 0.104 0.130 0.036 0.002 0.836 0.884 0.911 0.303 0.097
12 0.013 0.091 0.112 0.032 0.002 0.018 0.109 0.135 0.037 0.002 0.875 0.912 0.932 0.313 0.101
Avg 0.014 0.096 0.119 0.034 0.002 0.017 0.105 0.131 0.036 0.002 0.758 0.828 0.864 0.284 0.087
$M$-TSFLib 6 0.011 0.081 0.103 0.028 0.001 0.014 0.094 0.118 0.033 0.002 0.209 0.420 0.450 0.143 0.024
8 0.011 0.085 0.106 0.029 0.001 0.015 0.099 0.123 0.034 0.002 0.214 0.424 0.456 0.145 0.024
10 0.012 0.090 0.110 0.031 0.001 0.013 0.091 0.115 0.031 0.002 0.277 0.463 0.522 0.158 0.032
12 0.012 0.090 0.110 0.031 0.001 0.016 0.099 0.124 0.034 0.002 0.299 0.526 0.540 0.180 0.034
Avg 0.011 0.086 0.107 0.030 0.001 0.014 0.096 0.120 0.033 0.002 0.250 0.458 0.492 0.156 0.029
TaTS (ours) 6 0.008 0.077 0.090 0.027 0.001 0.009 0.080 0.097 0.028 0.001 0.140 0.312 0.367 0.106 0.016
8 0.008 0.077 0.090 0.027 0.001 0.008 0.078 0.091 0.027 0.001 0.212 0.426 0.454 0.145 0.024
10 0.009 0.079 0.093 0.027 0.001 0.009 0.079 0.092 0.027 0.001 0.302 0.510 0.544 0.174 0.034
12 0.008 0.076 0.091 0.026 0.001 0.009 0.080 0.096 0.028 0.001 0.222 0.428 0.466 0.145 0.025
Avg 0.008 0.077 0.091 0.027 0.001 0.009 0.079 0.094 0.028 0.001 0.219 0.419 0.458 0.142 0.025
Table 20: Full forecasting results for the Energy, Environment, and Health datasets using iTransformer, PatchTST, and Crossformer as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models iTransformer PatchTST Crossformer
(2024c) (2023) (2023)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Energy Uni-modal 12 0.112 0.231 0.305 0.940 10.877 0.105 0.229 0.301 1.120 40.016 0.138 0.262 0.338 1.121 31.808
24 0.222 0.352 0.438 1.626 35.711 0.241 0.363 0.458 1.507 41.679 0.281 0.399 0.482 1.619 46.008
36 0.306 0.409 0.511 1.767 58.611 0.304 0.408 0.512 1.782 60.280 0.331 0.443 0.541 3.296 229.673
48 0.435 0.509 0.617 2.454 95.250 0.427 0.502 0.610 2.405 91.255 0.422 0.521 0.626 4.574 347.533
Avg 0.269 0.375 0.468 1.697 50.112 0.269 0.376 0.470 1.704 58.307 0.293 0.406 0.497 2.652 163.756
$M$-TSFLib 12 0.107 0.228 0.300 0.999 17.267 0.115 0.242 0.311 1.357 30.739 0.126 0.251 0.326 1.141 30.653
24 0.223 0.354 0.443 1.692 45.580 0.236 0.361 0.453 1.506 35.353 0.268 0.389 0.477 1.875 69.289
36 0.313 0.424 0.520 1.916 63.383 0.311 0.413 0.518 1.885 64.280 0.338 0.454 0.553 2.956 202.647
48 0.423 0.507 0.615 2.494 101.822 0.426 0.501 0.610 2.404 95.489 0.433 0.534 0.637 4.716 415.767
Avg 0.267 0.378 0.469 1.775 57.013 0.272 0.379 0.473 1.788 56.465 0.291 0.407 0.498 2.672 179.589
TaTS (ours) 12 0.106 0.234 0.302 1.116 30.716 0.106 0.234 0.300 1.024 17.615 0.127 0.254 0.326 1.053 18.277
24 0.226 0.355 0.439 1.555 34.060 0.206 0.336 0.417 1.491 35.509 0.253 0.376 0.462 1.823 62.566
36 0.306 0.411 0.512 1.790 57.729 0.305 0.412 0.506 1.781 57.481 0.312 0.431 0.524 2.602 144.944
48 0.421 0.502 0.612 2.370 91.346 0.416 0.501 0.607 2.427 93.658 0.425 0.516 0.619 3.314 183.016
Avg 0.265 0.376 0.466 1.708 53.463 0.258 0.371 0.457 1.681 51.066 0.279 0.394 0.483 2.198 102.201
Environment Uni-modal 48 0.415 0.473 0.604 2.218 214.087 0.492 0.500 0.644 2.315 238.477 0.495 0.509 0.651 2.275 235.830
96 0.439 0.493 0.630 2.338 250.510 0.541 0.538 0.693 2.390 256.244 0.562 0.585 0.712 1.880 130.832
192 0.454 0.506 0.661 2.462 259.311 0.581 0.556 0.741 2.727 356.251 0.567 0.607 0.739 1.659 90.234
336 0.455 0.505 0.671 2.446 276.348 0.593 0.553 0.763 2.898 424.764 0.582 0.621 0.761 1.656 80.409
Avg 0.441 0.494 0.641 2.366 250.064 0.552 0.537 0.710 2.583 318.934 0.551 0.581 0.716 1.867 134.326
$M$-TSFLib 48 0.413 0.470 0.600 2.187 211.597 0.441 0.489 0.621 1.954 159.971 0.421 0.472 0.604 1.982 161.068
96 0.420 0.478 0.616 2.233 213.865 0.461 0.495 0.643 2.063 177.553 0.425 0.485 0.617 1.839 128.131
192 0.423 0.482 0.636 2.320 225.635 0.462 0.508 0.664 2.203 214.665 0.427 0.495 0.638 1.830 125.150
336 0.429 0.483 0.651 2.308 236.965 0.472 0.513 0.683 2.210 220.009 0.434 0.500 0.656 1.867 128.536
Avg 0.421 0.478 0.626 2.262 222.016 0.459 0.501 0.653 2.107 193.049 0.427 0.488 0.629 1.879 135.721
TaTS (ours) 48 0.268 0.370 0.480 1.140 21.157 0.271 0.377 0.483 1.115 19.355 0.274 0.373 0.485 1.152 20.952
96 0.267 0.370 0.488 1.123 20.955 0.279 0.376 0.498 1.176 21.868 0.284 0.391 0.505 1.127 18.484
192 0.272 0.366 0.508 1.215 23.115 0.272 0.366 0.508 1.215 23.342 0.283 0.415 0.523 1.039 14.150
336 0.261 0.369 0.508 1.174 22.437 0.269 0.366 0.516 1.222 23.081 0.294 0.431 0.541 1.031 12.734
Avg 0.267 0.369 0.496 1.163 21.916 0.273 0.371 0.501 1.182 21.912 0.284 0.403 0.513 1.087 16.580
Health Uni-modal 12 1.171 0.674 0.967 2.596 149.272 1.267 0.735 0.999 3.651 262.953 1.453 0.809 1.086 2.699 145.979
24 1.594 0.807 1.169 2.832 136.133 1.681 0.846 1.167 3.371 160.263 1.537 0.825 1.153 2.744 100.232
36 1.742 0.862 1.253 2.893 118.496 1.819 0.918 1.281 3.751 243.238 1.565 0.831 1.173 2.687 108.731
48 1.840 0.923 1.313 3.263 153.559 1.842 0.921 1.310 3.192 152.876 1.586 0.841 1.208 2.730 165.914
Avg 1.587 0.817 1.175 2.896 139.365 1.652 0.855 1.189 3.491 204.832 1.535 0.827 1.155 2.715 130.214
$M$-TSFLib 12 0.987 0.696 0.928 3.334 190.818 1.029 0.710 0.945 3.400 182.961 1.017 0.664 0.933 2.796 171.177
24 1.388 0.787 1.113 3.457 169.217 1.288 0.769 1.053 3.207 129.039 1.318 0.747 1.072 2.917 161.566
36 1.672 0.877 1.216 3.674 176.904 1.460 0.831 1.152 3.492 168.807 1.357 0.775 1.113 3.210 192.422
48 1.737 0.902 1.270 3.531 155.800 1.612 0.878 1.231 3.317 127.827 1.399 0.788 1.146 2.821 149.248
Avg 1.446 0.816 1.132 3.499 173.185 1.347 0.797 1.095 3.354 152.159 1.273 0.744 1.066 2.936 168.603
TaTS (ours) 12 0.939 0.649 0.892 2.764 138.387 0.990 0.659 0.912 2.721 151.988 0.974 0.661 0.911 2.838 163.149
24 1.251 0.712 1.032 2.744 112.080 1.288 0.764 1.050 3.364 155.937 1.254 0.741 1.062 3.104 170.871
36 1.489 0.781 1.147 2.885 112.555 1.397 0.780 1.122 2.638 93.114 1.306 0.741 1.083 2.805 152.713
48 1.581 0.834 1.215 2.909 107.995 1.456 0.808 1.165 2.623 88.216 1.369 0.770 1.136 2.770 144.765
Avg 1.315 0.744 1.071 2.825 117.754 1.283 0.753 1.062 2.837 122.314 1.226 0.728 1.048 2.879 157.874
Table 21: Full forecasting results for the Security, Social Good, and Traffic datasets using iTransformer, PatchTST, and Crossformer as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models iTransformer PatchTST Crossformer
(2024c) (2023) (2023)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Security Uni-modal 6 113.573 5.698 10.657 6.008 1020.167 108.795 5.119 10.430 4.255 549.239 124.972 6.077 11.179 2.435 113.813
8 115.878 5.743 10.765 5.977 1035.504 113.534 5.524 10.655 3.681 424.901 126.513 6.237 11.248 1.909 67.155
10 116.950 5.615 10.814 3.109 332.477 114.820 5.484 10.715 2.523 185.108 127.701 6.346 11.300 1.678 70.642
12 117.396 5.585 10.835 2.645 262.692 114.253 5.357 10.689 1.951 117.032 128.686 6.449 11.344 1.484 35.512
Avg 115.949 5.660 10.768 4.435 662.710 112.850 5.371 10.622 3.103 319.070 126.968 6.277 11.268 1.877 71.781
$M$-TSFLib 6 115.170 5.403 10.732 4.284 482.301 109.240 5.189 10.452 4.386 560.668 123.515 5.958 11.114 2.771 175.140
8 116.158 5.544 10.778 5.742 985.203 113.248 5.445 10.642 3.237 322.665 124.714 6.100 11.168 2.169 126.854
10 117.340 5.633 10.832 3.090 320.102 114.109 5.442 10.682 2.500 196.618 127.701 6.346 11.300 1.678 70.642
12 116.694 5.548 10.803 2.560 243.369 114.764 5.398 10.713 2.136 165.771 126.985 6.327 11.269 1.510 42.745
Avg 116.341 5.532 10.786 3.919 507.744 112.840 5.369 10.622 3.065 311.430 125.729 6.183 11.213 2.032 103.845
TaTS (ours) 6 107.113 4.856 10.350 3.600 320.907 106.160 4.696 10.303 3.316 282.430 122.887 5.915 11.085 2.831 191.546
8 112.560 5.204 10.609 3.258 345.121 108.803 5.052 10.431 3.082 312.449 124.302 6.067 11.149 2.656 178.725
10 113.789 5.227 10.667 2.488 204.306 111.110 5.090 10.541 2.423 205.196 126.203 6.258 11.234 1.720 60.667
12 114.754 5.318 10.712 2.057 140.148 112.699 5.237 10.616 2.185 171.929 127.263 6.352 11.281 1.433 36.934
Avg 112.054 5.151 10.584 2.851 252.621 109.693 5.019 10.473 2.752 243.001 125.164 6.148 11.187 2.160 116.968
Social Good Uni-modal 6 1.129 0.438 0.741 1.395 72.570 1.047 0.443 0.749 1.420 70.855 0.791 0.431 0.681 1.320 39.179
8 1.133 0.466 0.770 1.448 67.692 1.102 0.458 0.741 1.502 68.611 0.832 0.414 0.668 0.917 13.263
10 1.275 0.496 0.808 1.718 112.603 1.089 0.469 0.749 1.568 75.225 0.916 0.463 0.728 0.709 6.748
12 1.313 0.534 0.841 2.052 159.814 1.148 0.610 0.866 2.185 94.833 0.921 0.559 0.766 1.682 54.438
Avg 1.212 0.483 0.790 1.653 103.170 1.097 0.495 0.776 1.669 77.381 0.865 0.467 0.711 1.157 28.407
$M$-TSFLib 6 1.061 0.451 0.752 1.399 60.240 1.023 0.442 0.741 1.333 62.404 0.753 0.350 0.604 0.753 8.426
8 1.175 0.498 0.789 1.390 59.283 1.111 0.533 0.807 1.295 34.722 0.814 0.357 0.618 0.576 5.614
10 1.253 0.561 0.863 1.627 66.938 1.049 0.520 0.799 1.504 54.694 0.866 0.492 0.726 1.127 10.642
12 1.298 0.570 0.872 1.710 72.310 1.110 0.566 0.840 1.632 50.501 0.917 0.395 0.650 0.693 3.899
Avg 1.197 0.520 0.819 1.531 64.693 1.073 0.515 0.797 1.441 50.580 0.837 0.398 0.649 0.787 7.145
TaTS (ours) 6 0.942 0.398 0.677 1.221 49.723 0.923 0.436 0.722 1.153 18.695 0.711 0.407 0.631 0.749 6.772
8 0.967 0.433 0.713 1.430 50.480 0.900 0.461 0.713 1.279 22.919 0.748 0.453 0.646 0.963 7.987
10 0.994 0.463 0.740 1.538 54.466 0.996 0.461 0.728 1.348 42.272 0.800 0.373 0.615 0.724 5.977
12 1.045 0.514 0.780 1.753 61.884 1.069 0.501 0.770 1.549 60.142 0.857 0.415 0.663 0.827 7.195
Avg 0.987 0.452 0.728 1.486 54.138 0.972 0.465 0.733 1.332 36.007 0.779 0.412 0.639 0.816 6.983
Traffic Uni-modal 6 0.203 0.228 0.393 0.216 0.307 0.182 0.252 0.377 0.285 0.513 0.227 0.394 0.472 0.340 0.388
8 0.209 0.236 0.399 0.224 0.321 0.167 0.226 0.348 0.255 0.450 0.216 0.382 0.458 0.324 0.333
10 0.211 0.243 0.398 0.229 0.321 0.178 0.242 0.362 0.270 0.464 0.202 0.363 0.441 0.313 0.327
12 0.231 0.246 0.392 0.281 0.532 0.226 0.250 0.393 0.300 0.580 0.212 0.365 0.450 0.337 0.399
Avg 0.213 0.238 0.395 0.238 0.370 0.188 0.242 0.370 0.278 0.502 0.214 0.376 0.455 0.329 0.362
$M$-TSFLib 6 0.187 0.338 0.422 0.292 0.324 0.165 0.229 0.353 0.248 0.391 0.184 0.335 0.418 0.307 0.371
8 0.197 0.355 0.433 0.297 0.303 0.163 0.218 0.346 0.237 0.373 0.183 0.331 0.416 0.305 0.367
10 0.190 0.338 0.422 0.285 0.283 0.174 0.235 0.357 0.252 0.402 0.184 0.331 0.416 0.303 0.363
12 0.222 0.358 0.459 0.333 0.410 0.211 0.237 0.378 0.281 0.524 0.200 0.340 0.431 0.329 0.427
Avg 0.199 0.347 0.434 0.302 0.330 0.178 0.230 0.359 0.255 0.422 0.188 0.334 0.420 0.311 0.382
TaTS (ours) 6 0.174 0.218 0.351 0.227 0.348 0.155 0.204 0.334 0.225 0.359 0.159 0.275 0.372 0.273 0.392
8 0.177 0.213 0.357 0.215 0.319 0.162 0.210 0.334 0.235 0.389 0.166 0.294 0.385 0.279 0.369
10 0.186 0.225 0.366 0.223 0.324 0.167 0.214 0.340 0.233 0.368 0.163 0.285 0.378 0.277 0.386
12 0.213 0.212 0.361 0.241 0.451 0.204 0.209 0.356 0.249 0.468 0.184 0.291 0.395 0.296 0.428
Avg 0.187 0.217 0.359 0.227 0.361 0.172 0.209 0.341 0.235 0.396 0.168 0.286 0.383 0.281 0.394
Table 22: Full forecasting results for the Agriculture, Climate, and Economy datasets using DLinear, FEDformer, and FiLM as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models DLinear FEDformer FiLM
(2023) (2022b) (2022a)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Agriculture Uni-modal 6 0.170 0.312 0.395 0.135 0.028 0.091 0.239 0.301 0.110 0.019 0.088 0.207 0.290 0.092 0.015
8 0.195 0.340 0.423 0.145 0.031 0.127 0.281 0.355 0.127 0.024 0.115 0.230 0.318 0.097 0.017
10 0.218 0.355 0.448 0.150 0.034 0.154 0.313 0.392 0.142 0.030 0.151 0.255 0.351 0.103 0.020
12 0.308 0.407 0.514 0.165 0.041 0.180 0.313 0.415 0.134 0.028 0.200 0.333 0.433 0.140 0.030
Avg 0.223 0.354 0.445 0.149 0.034 0.138 0.286 0.366 0.128 0.025 0.139 0.256 0.348 0.108 0.021
$M$-TSFLib 6 0.166 0.314 0.396 0.137 0.028 0.079 0.219 0.280 0.102 0.016 0.089 0.208 0.291 0.092 0.015
8 0.193 0.342 0.425 0.148 0.032 0.112 0.269 0.335 0.125 0.024 0.112 0.225 0.314 0.095 0.016
10 0.214 0.358 0.450 0.153 0.035 0.153 0.292 0.387 0.128 0.026 0.156 0.262 0.357 0.106 0.020
12 0.300 0.408 0.514 0.167 0.042 0.180 0.319 0.416 0.137 0.028 0.204 0.338 0.437 0.142 0.031
Avg 0.218 0.355 0.446 0.151 0.034 0.131 0.275 0.354 0.123 0.024 0.140 0.258 0.350 0.109 0.021
TaTS (ours) 6 0.164 0.311 0.392 0.136 0.028 0.082 0.222 0.286 0.102 0.016 0.087 0.205 0.289 0.091 0.015
8 0.192 0.343 0.425 0.148 0.032 0.111 0.256 0.330 0.115 0.020 0.110 0.223 0.312 0.094 0.016
10 0.215 0.358 0.451 0.152 0.035 0.147 0.288 0.377 0.126 0.025 0.146 0.249 0.345 0.100 0.019
12 0.287 0.392 0.496 0.159 0.038 0.183 0.339 0.418 0.145 0.029 0.196 0.328 0.429 0.138 0.029
Avg 0.214 0.351 0.441 0.149 0.033 0.131 0.276 0.353 0.122 0.023 0.135 0.251 0.344 0.106 0.020
Climate Uni-modal 6 1.158 0.866 1.076 3.462 204.312 1.206 0.909 1.098 4.551 394.050 1.277 0.912 1.129 4.953 410.091
8 1.191 0.874 1.091 3.400 190.434 1.175 0.897 1.084 3.680 176.984 1.158 0.862 1.076 3.282 162.026
10 1.225 0.882 1.107 3.392 185.912 1.199 0.893 1.095 3.215 134.937 1.160 0.857 1.077 2.152 48.217
12 1.185 0.868 1.089 2.856 126.200 1.190 0.875 1.091 2.883 99.696 1.487 1.012 1.219 3.993 164.875
Avg 1.190 0.872 1.091 3.277 176.715 1.192 0.893 1.092 3.582 201.417 1.270 0.911 1.125 3.595 196.302
$M$-TSFLib 6 1.074 0.822 1.036 2.997 154.862 1.012 0.795 1.006 2.492 59.880 1.165 0.869 1.079 3.798 270.733
8 1.108 0.844 1.053 3.117 160.555 1.003 0.787 1.001 2.265 34.593 1.173 0.867 1.083 3.599 191.009
10 1.123 0.844 1.059 3.233 167.366 1.011 0.810 1.004 2.133 24.604 1.134 0.843 1.065 1.946 31.959
12 1.112 0.838 1.054 2.815 116.283 1.019 0.795 1.007 2.321 40.325 1.245 0.904 1.116 3.641 223.644
Avg 1.104 0.837 1.050 3.040 149.767 1.011 0.797 1.004 2.303 39.851 1.179 0.871 1.086 3.246 179.336
TaTS (ours) 6 0.905 0.749 0.951 2.025 36.745 0.893 0.748 0.945 1.806 13.948 0.912 0.758 0.955 2.598 82.597
8 0.926 0.756 0.962 2.018 37.680 0.937 0.771 0.968 1.893 16.622 0.917 0.751 0.957 1.939 26.990
10 0.943 0.764 0.971 1.989 34.385 0.924 0.751 0.960 1.756 14.796 0.947 0.759 0.972 1.597 10.958
12 0.950 0.766 0.973 1.938 30.964 0.952 0.770 0.974 1.921 21.544 1.005 0.821 1.003 2.686 68.310
Avg 0.931 0.759 0.964 1.992 34.944 0.926 0.760 0.962 1.844 16.727 0.945 0.772 0.972 2.205 47.214
Economy Uni-modal 6 0.056 0.189 0.236 0.065 0.006 0.042 0.168 0.203 0.058 0.005 0.021 0.115 0.146 0.039 0.002
8 0.056 0.191 0.237 0.066 0.006 0.039 0.162 0.197 0.056 0.005 0.028 0.132 0.168 0.045 0.003
10 0.047 0.173 0.216 0.060 0.006 0.036 0.153 0.188 0.053 0.004 0.026 0.134 0.162 0.045 0.003
12 0.075 0.216 0.272 0.074 0.009 0.053 0.183 0.225 0.063 0.006 0.027 0.133 0.163 0.046 0.003
Avg 0.058 0.192 0.240 0.066 0.007 0.042 0.166 0.203 0.058 0.005 0.025 0.129 0.160 0.044 0.003
$M$-TSFLib 6 0.059 0.200 0.243 0.070 0.007 0.035 0.153 0.188 0.053 0.004 0.018 0.104 0.133 0.036 0.002
8 0.062 0.194 0.248 0.068 0.007 0.043 0.170 0.206 0.059 0.005 0.031 0.138 0.177 0.047 0.003
10 0.064 0.195 0.253 0.068 0.008 0.040 0.160 0.200 0.056 0.005 0.026 0.133 0.160 0.045 0.003
12 0.049 0.180 0.221 0.062 0.006 0.024 0.129 0.157 0.045 0.003 0.029 0.140 0.170 0.048 0.003
Avg 0.058 0.192 0.241 0.067 0.007 0.035 0.153 0.188 0.053 0.004 0.026 0.129 0.160 0.044 0.003
TaTS (ours) 6 0.020 0.115 0.141 0.040 0.002 0.012 0.093 0.111 0.033 0.002 0.009 0.080 0.096 0.028 0.001
8 0.020 0.115 0.142 0.040 0.002 0.014 0.099 0.120 0.034 0.002 0.009 0.079 0.096 0.028 0.001
10 0.019 0.112 0.137 0.039 0.002 0.016 0.106 0.126 0.037 0.002 0.009 0.079 0.096 0.027 0.001
12 0.025 0.127 0.157 0.044 0.003 0.017 0.107 0.129 0.037 0.002 0.009 0.081 0.096 0.028 0.001
Avg 0.021 0.117 0.144 0.041 0.002 0.015 0.101 0.121 0.035 0.002 0.009 0.080 0.096 0.028 0.001
Table 23: Full forecasting results for the Energy, Environment, and Health datasets using DLinear, FEDformer, and FiLM as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models DLinear FEDformer FiLM
(2023) (2022b) (2022a)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Energy Uni-modal 12 0.136 0.264 0.335 0.947 10.373 0.095 0.212 0.281 0.895 9.380 0.118 0.245 0.314 1.208 25.502
24 0.261 0.385 0.467 1.453 33.690 0.170 0.303 0.385 1.573 47.676 0.221 0.347 0.433 1.414 29.039
36 0.335 0.437 0.528 1.780 56.178 0.249 0.373 0.475 2.165 133.305 0.335 0.434 0.533 1.852 60.033
48 0.431 0.498 0.603 2.294 83.886 0.445 0.514 0.643 4.141 340.071 0.437 0.512 0.620 2.414 89.205
Avg 0.291 0.396 0.483 1.619 46.032 0.240 0.351 0.446 2.194 132.608 0.278 0.385 0.475 1.722 50.945
$M$-TSFLib 12 0.133 0.262 0.332 0.950 10.725 0.098 0.227 0.289 1.083 15.260 0.119 0.247 0.315 1.202 24.419
24 0.256 0.380 0.461 1.428 32.596 0.172 0.300 0.388 1.425 47.727 0.224 0.349 0.436 1.415 28.417
36 0.340 0.442 0.533 1.803 56.982 0.252 0.376 0.477 2.120 124.465 0.332 0.431 0.531 1.839 59.454
48 0.428 0.496 0.601 2.287 83.443 0.430 0.511 0.611 2.775 146.406 0.440 0.514 0.623 2.416 89.684
Avg 0.289 0.395 0.482 1.617 45.936 0.238 0.354 0.441 1.851 83.465 0.279 0.385 0.476 1.718 50.493
TaTS (ours) 12 0.132 0.260 0.330 0.954 11.225 0.090 0.210 0.275 0.880 12.877 0.118 0.245 0.314 1.202 25.174
24 0.236 0.359 0.441 1.359 29.647 0.172 0.305 0.387 1.375 44.042 0.222 0.347 0.434 1.410 28.640
36 0.340 0.442 0.533 1.805 57.123 0.250 0.372 0.472 2.412 152.144 0.311 0.414 0.514 1.773 55.535
48 0.425 0.493 0.597 2.285 84.710 0.435 0.531 0.631 3.501 222.058 0.434 0.510 0.618 2.399 88.602
Avg 0.283 0.388 0.475 1.601 45.676 0.237 0.355 0.441 2.042 107.780 0.271 0.379 0.470 1.696 49.488
Environment Uni-modal 48 0.478 0.531 0.646 1.791 119.315 0.505 0.543 0.671 1.932 128.444 0.494 0.502 0.644 2.323 241.494
96 0.562 0.608 0.724 1.539 68.485 0.465 0.524 0.657 2.440 237.225 0.581 0.546 0.707 2.639 314.062
192 0.592 0.608 0.748 1.938 132.178 0.510 0.556 0.702 2.656 297.619 0.612 0.559 0.759 3.081 461.654
336 0.600 0.618 0.768 1.938 134.420 0.531 0.573 0.725 2.528 300.421 0.621 0.566 0.779 2.989 440.819
Avg 0.558 0.591 0.722 1.801 113.600 0.503 0.549 0.689 2.389 240.927 0.577 0.543 0.722 2.758 364.507
$M$-TSFLib 48 0.414 0.474 0.598 1.888 142.797 0.413 0.468 0.599 1.867 128.047 0.469 0.484 0.652 2.046 177.399
96 0.420 0.489 0.615 1.743 112.990 0.413 0.479 0.613 1.998 138.883 0.476 0.493 0.662 1.987 173.914
192 0.439 0.521 0.650 1.659 92.080 0.423 0.489 0.635 2.205 197.315 0.477 0.486 0.674 2.135 218.811
336 0.444 0.526 0.664 1.680 95.029 0.445 0.507 0.665 1.850 125.600 0.490 0.497 0.696 2.233 243.544
Avg 0.429 0.502 0.632 1.742 110.724 0.423 0.486 0.628 1.980 147.461 0.478 0.490 0.671 2.100 203.417
TaTS (ours) 48 0.272 0.385 0.486 1.083 17.610 0.272 0.369 0.484 1.172 21.826 0.269 0.373 0.481 1.124 19.765
96 0.301 0.427 0.528 1.034 13.731 0.271 0.371 0.495 1.143 22.305 0.279 0.377 0.499 1.180 21.986
192 0.306 0.443 0.544 1.007 11.494 0.277 0.379 0.501 1.145 22.438 0.271 0.367 0.507 1.202 22.318
336 0.313 0.456 0.557 0.996 10.408 0.278 0.393 0.524 1.137 18.379 0.267 0.366 0.514 1.210 22.394
Avg 0.298 0.428 0.529 1.030 13.311 0.275 0.378 0.501 1.149 21.237 0.272 0.371 0.500 1.179 21.616
Health Uni-modal 12 1.595 0.808 1.113 2.599 252.012 1.051 0.756 0.975 4.069 371.454 1.900 1.008 1.278 6.208 1054.818
24 1.778 0.835 1.170 2.742 312.129 1.493 0.933 1.176 4.818 434.785 1.946 0.973 1.287 4.194 322.873
36 1.759 0.850 1.241 2.738 263.854 1.661 0.980 1.252 5.182 426.708 2.029 1.013 1.352 4.495 346.381
48 1.818 0.899 1.299 2.964 278.936 1.737 0.969 1.285 4.775 392.270 2.054 1.026 1.379 4.356 305.131
Avg 1.737 0.848 1.206 2.761 276.733 1.486 0.909 1.172 4.711 406.304 1.982 1.005 1.324 4.813 507.301
$M$-TSFLib 12 1.403 0.763 1.061 2.065 64.534 0.978 0.685 0.930 3.371 273.283 1.406 0.915 1.141 4.787 421.324
24 1.544 0.782 1.114 2.167 68.258 1.272 0.804 1.073 3.538 212.521 1.730 0.975 1.257 4.258 251.992
36 1.600 0.824 1.185 2.264 56.729 1.344 0.828 1.121 3.768 222.549 1.769 0.944 1.267 3.785 205.906
48 1.617 0.830 1.222 2.368 87.198 1.416 0.852 1.163 3.913 243.392 1.793 0.960 1.286 3.626 178.720
Avg 1.541 0.800 1.145 2.216 69.180 1.252 0.792 1.072 3.647 237.936 1.675 0.949 1.238 4.114 264.486
TaTS (ours) 12 1.273 0.742 1.027 2.702 159.975 0.966 0.671 0.925 3.309 294.367 1.211 0.794 1.035 3.570 236.037
24 1.421 0.788 1.090 3.221 254.155 1.264 0.823 1.080 4.068 289.731 1.414 0.827 1.107 3.160 130.456
36 1.449 0.787 1.133 2.812 148.843 1.320 0.801 1.110 3.885 280.248 1.497 0.851 1.165 3.274 133.041
48 1.505 0.832 1.183 2.990 186.495 1.426 0.868 1.164 4.140 264.871 1.562 0.878 1.209 3.225 121.440
Avg 1.412 0.787 1.108 2.931 187.367 1.244 0.791 1.070 3.851 282.304 1.421 0.838 1.129 3.307 155.243
Table 24: Full forecasting results for the Security, Social Good, and Traffic datasets using DLinear, FEDformer, and FiLM as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models DLinear FEDformer FiLM
(2023) (2022b) (2022a)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Security Uni-modal 6 107.046 4.531 10.346 3.489 324.650 112.106 4.830 10.588 4.456 571.690 115.130 5.587 10.730 5.029 637.394
8 107.700 4.649 10.378 2.992 295.719 113.897 5.200 10.672 3.223 355.929 112.097 5.034 10.588 3.228 336.503
10 109.360 4.759 10.458 2.449 216.866 117.484 5.467 10.839 2.563 216.459 112.267 4.901 10.596 2.310 188.201
12 112.343 4.904 10.599 1.980 149.507 114.452 5.135 10.698 2.697 329.415 122.711 6.425 11.077 3.554 437.883
Avg 109.112 4.711 10.445 2.728 246.685 114.485 5.158 10.699 3.235 368.373 115.551 5.487 10.748 3.530 399.995
$M$-TSFLib 6 106.121 4.545 10.301 3.788 386.715 109.854 4.690 10.481 4.252 541.638 105.809 4.755 10.286 4.286 525.309
8 107.445 4.670 10.366 3.101 319.555 114.833 5.233 10.716 3.119 330.374 108.045 4.638 10.394 2.727 249.934
10 108.713 4.758 10.427 2.613 251.113 115.261 5.294 10.736 2.563 233.233 111.391 4.843 10.554 1.953 137.888
12 109.848 4.876 10.481 2.373 227.133 115.001 5.211 10.724 2.558 291.461 111.541 5.354 10.561 3.023 363.599
Avg 108.032 4.712 10.394 2.969 296.129 113.737 5.107 10.664 3.123 349.177 109.197 4.897 10.449 2.997 319.183
TaTS (ours) 6 106.015 4.516 10.296 3.850 404.799 106.015 4.532 10.296 4.542 597.429 105.482 4.511 10.270 4.024 451.366
8 107.477 4.623 10.367 3.050 312.447 107.678 4.618 10.377 3.303 371.986 107.657 4.792 10.376 3.150 310.920
10 108.505 4.728 10.417 2.632 260.096 107.301 4.885 10.359 3.138 375.009 109.886 4.771 10.483 2.338 193.865
12 109.717 4.836 10.475 2.336 224.134 108.506 4.837 10.417 2.738 331.710 108.375 4.870 10.410 2.549 267.339
Avg 107.928 4.676 10.389 2.967 300.369 107.375 4.718 10.362 3.430 419.034 107.850 4.736 10.385 3.015 305.873
Social Good Uni-modal 6 1.018 0.627 0.859 1.503 19.120 0.821 0.386 0.649 1.086 28.066 1.123 0.604 0.878 2.432 149.647
8 1.137 0.702 0.929 1.607 19.248 0.929 0.451 0.726 1.061 17.701 1.116 0.542 0.841 2.275 128.188
10 1.210 0.755 0.972 1.732 21.287 1.059 0.478 0.784 1.469 52.985 1.154 0.601 0.882 1.823 42.492
12 1.238 0.763 0.982 1.705 20.522 1.105 0.588 0.846 1.876 76.675 1.653 0.869 1.125 2.998 140.138
Avg 1.151 0.712 0.935 1.637 20.044 0.979 0.476 0.751 1.373 43.857 1.261 0.654 0.931 2.382 115.116
$M$-TSFLib 6 0.946 0.583 0.810 1.481 21.957 0.816 0.395 0.667 1.182 33.446 1.070 0.529 0.824 2.269 159.103
8 1.095 0.678 0.904 1.596 20.076 0.910 0.438 0.717 1.010 15.845 1.120 0.584 0.869 2.138 83.150
10 1.128 0.709 0.924 1.713 22.953 1.032 0.456 0.757 1.395 53.857 1.128 0.539 0.839 2.235 120.886
12 1.164 0.724 0.940 1.655 20.260 1.091 0.560 0.838 1.737 71.467 1.624 0.852 1.111 3.005 149.026
Avg 1.083 0.673 0.894 1.611 21.312 0.962 0.462 0.745 1.331 43.654 1.236 0.626 0.911 2.412 128.041
TaTS (ours) 6 0.859 0.516 0.740 1.453 29.534 0.746 0.365 0.620 0.983 23.097 0.992 0.588 0.834 1.623 30.115
8 0.963 0.592 0.812 1.582 29.073 0.880 0.448 0.701 1.221 34.945 1.001 0.550 0.806 1.609 28.399
10 1.118 0.703 0.917 1.711 23.312 0.944 0.453 0.720 1.138 28.381 1.065 0.583 0.837 1.447 17.661
12 1.085 0.676 0.888 1.657 24.350 0.983 0.453 0.718 1.315 51.258 1.358 0.782 1.025 1.920 27.141
Avg 1.006 0.622 0.839 1.601 26.567 0.888 0.430 0.690 1.164 34.420 1.104 0.626 0.876 1.650 25.829
Traffic Uni-modal 6 0.221 0.358 0.461 0.357 0.549 0.202 0.294 0.416 0.285 0.377 0.202 0.302 0.426 0.346 0.682
8 0.220 0.357 0.458 0.355 0.552 0.184 0.234 0.364 0.259 0.461 0.203 0.332 0.436 0.341 0.564
10 0.217 0.347 0.451 0.358 0.601 0.198 0.251 0.386 0.271 0.474 0.203 0.340 0.438 0.323 0.447
12 0.261 0.374 0.499 0.395 0.663 0.237 0.275 0.410 0.315 0.600 0.250 0.280 0.426 0.347 0.719
Avg 0.230 0.359 0.467 0.366 0.591 0.205 0.264 0.394 0.283 0.478 0.215 0.314 0.431 0.339 0.603
$M$-TSFLib 6 0.204 0.332 0.439 0.321 0.465 0.180 0.245 0.372 0.258 0.408 0.197 0.294 0.419 0.339 0.664
8 0.202 0.335 0.435 0.320 0.448 0.178 0.225 0.352 0.253 0.458 0.195 0.322 0.427 0.332 0.540
10 0.205 0.325 0.435 0.320 0.478 0.184 0.234 0.364 0.258 0.446 0.195 0.328 0.428 0.312 0.424
12 0.225 0.329 0.451 0.351 0.602 0.228 0.247 0.388 0.292 0.570 0.240 0.258 0.403 0.315 0.631
Avg 0.209 0.330 0.440 0.328 0.498 0.193 0.238 0.369 0.265 0.471 0.207 0.300 0.419 0.325 0.565
TaTS (ours) 6 0.184 0.312 0.412 0.301 0.435 0.159 0.208 0.329 0.228 0.355 0.157 0.228 0.351 0.254 0.426
8 0.185 0.302 0.409 0.308 0.500 0.159 0.209 0.329 0.230 0.368 0.162 0.260 0.370 0.266 0.401
10 0.184 0.297 0.403 0.309 0.524 0.160 0.206 0.323 0.229 0.370 0.169 0.269 0.380 0.262 0.358
12 0.199 0.291 0.413 0.309 0.489 0.213 0.224 0.368 0.269 0.518 0.215 0.236 0.383 0.298 0.612
Avg 0.188 0.300 0.409 0.307 0.487 0.173 0.212 0.337 0.239 0.403 0.176 0.248 0.371 0.270 0.449
Table 25: Full forecasting results for the Agriculture, Climate, and Economy datasets using Autoformer, Informer, and Transformer as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models Autoformer Informer Transformer
(2021) (2021) (2017)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Agriculture Uni-modal 6 0.109 0.255 0.329 0.116 0.022 0.451 0.557 0.626 0.240 0.071 0.229 0.326 0.420 0.133 0.032
8 0.135 0.278 0.366 0.126 0.026 0.569 0.633 0.702 0.270 0.087 0.328 0.432 0.502 0.177 0.045
10 0.173 0.311 0.413 0.138 0.031 0.618 0.633 0.724 0.263 0.088 0.358 0.422 0.510 0.166 0.045
12 0.213 0.345 0.457 0.149 0.036 0.756 0.698 0.789 0.284 0.101 0.500 0.555 0.633 0.224 0.064
Avg 0.158 0.297 0.391 0.132 0.029 0.599 0.630 0.710 0.264 0.087 0.354 0.434 0.516 0.175 0.046
$M$-TSFLib 6 0.095 0.221 0.306 0.100 0.018 0.218 0.352 0.422 0.148 0.032 0.197 0.319 0.392 0.133 0.029
8 0.143 0.278 0.372 0.122 0.025 0.306 0.429 0.496 0.179 0.044 0.205 0.303 0.384 0.121 0.027
10 0.176 0.311 0.416 0.137 0.031 0.301 0.398 0.480 0.160 0.040 0.215 0.332 0.411 0.135 0.030
12 0.217 0.341 0.453 0.144 0.034 0.428 0.479 0.579 0.190 0.054 0.378 0.453 0.531 0.179 0.048
Avg 0.158 0.288 0.387 0.126 0.027 0.313 0.414 0.494 0.169 0.042 0.249 0.352 0.429 0.142 0.034
TaTS (ours) 6 0.076 0.205 0.269 0.092 0.014 0.162 0.265 0.342 0.107 0.022 0.150 0.348 0.387 0.166 0.034
8 0.101 0.234 0.315 0.105 0.019 0.213 0.311 0.392 0.124 0.028 0.192 0.298 0.374 0.120 0.025
10 0.138 0.299 0.371 0.134 0.026 0.296 0.388 0.465 0.154 0.038 0.209 0.296 0.384 0.114 0.026
12 0.186 0.324 0.425 0.141 0.030 0.349 0.426 0.506 0.167 0.044 0.213 0.310 0.401 0.121 0.025
Avg 0.125 0.266 0.345 0.118 0.022 0.255 0.348 0.426 0.138 0.033 0.191 0.313 0.387 0.130 0.028
Climate Uni-modal 6 1.116 0.861 1.056 4.456 519.418 1.084 0.829 1.041 1.617 15.451 1.032 0.817 1.016 1.656 13.521
8 1.117 0.856 1.057 3.329 177.957 1.087 0.829 1.043 2.464 66.852 1.135 0.856 1.065 1.868 22.687
10 1.170 0.886 1.081 3.059 106.074 1.117 0.849 1.056 2.071 49.757 1.103 0.844 1.049 2.355 81.661
12 1.123 0.855 1.059 3.364 125.779 1.150 0.859 1.071 2.557 94.732 1.100 0.841 1.048 1.663 24.078
Avg 1.131 0.865 1.063 3.552 232.307 1.110 0.841 1.053 2.177 56.698 1.092 0.839 1.044 1.885 35.487
$M$-TSFLib 6 1.021 0.817 1.010 2.871 82.634 1.019 0.811 1.009 2.844 134.946 0.962 0.771 0.981 2.030 23.607
8 1.039 0.819 1.018 3.018 99.283 0.970 0.772 0.985 2.210 43.152 0.991 0.783 0.995 1.976 22.312
10 1.074 0.838 1.035 2.783 92.522 1.006 0.794 1.002 2.311 67.013 1.026 0.790 1.010 2.755 87.765
12 1.078 0.836 1.038 2.731 78.154 1.009 0.791 0.999 2.499 86.993 1.011 0.789 1.000 1.895 27.637
Avg 1.053 0.827 1.025 2.851 88.148 1.001 0.792 0.999 2.466 83.026 0.998 0.783 0.996 2.164 40.330
TaTS (ours) 6 0.880 0.741 0.938 2.222 37.958 0.881 0.741 0.939 2.217 57.260 0.859 0.728 0.926 1.970 35.573
8 0.937 0.768 0.968 2.136 36.109 0.909 0.749 0.953 1.981 29.665 0.913 0.761 0.955 2.524 73.564
10 1.027 0.817 1.011 2.260 31.311 0.973 0.771 0.983 2.020 27.051 0.976 0.767 0.985 2.179 31.989
12 1.075 0.831 1.033 2.317 35.133 0.958 0.764 0.974 1.727 15.480 0.930 0.757 0.960 1.871 23.016
Avg 0.980 0.789 0.987 2.234 35.128 0.930 0.756 0.962 1.986 32.364 0.920 0.753 0.957 2.136 41.035
Economy Uni-modal 6 0.083 0.222 0.288 0.077 0.010 0.877 0.896 0.930 0.308 0.102 0.276 0.444 0.517 0.150 0.031
8 0.069 0.210 0.263 0.073 0.008 1.606 1.238 1.260 0.425 0.187 0.676 0.797 0.816 0.274 0.078
10 0.070 0.209 0.261 0.072 0.008 1.409 1.153 1.179 0.395 0.162 0.601 0.750 0.768 0.257 0.069
12 0.062 0.186 0.245 0.064 0.007 1.407 1.153 1.183 0.396 0.163 0.781 0.854 0.879 0.293 0.091
Avg 0.071 0.207 0.264 0.071 0.008 1.325 1.110 1.138 0.381 0.153 0.584 0.711 0.745 0.243 0.067
$M$-TSFLib 6 0.049 0.173 0.221 0.060 0.006 0.293 0.490 0.531 0.167 0.033 0.110 0.285 0.322 0.096 0.012
8 0.045 0.171 0.211 0.060 0.005 0.445 0.643 0.660 0.220 0.051 0.157 0.359 0.389 0.122 0.018
10 0.071 0.214 0.267 0.075 0.009 0.471 0.664 0.680 0.227 0.054 0.240 0.450 0.483 0.154 0.028
12 0.068 0.209 0.260 0.073 0.008 0.518 0.674 0.715 0.231 0.060 0.347 0.569 0.584 0.195 0.040
Avg 0.058 0.192 0.240 0.067 0.007 0.432 0.618 0.646 0.211 0.049 0.213 0.416 0.445 0.142 0.025
TaTS (ours) 6 0.021 0.115 0.144 0.040 0.002 0.247 0.472 0.488 0.161 0.028 0.029 0.137 0.169 0.047 0.003
8 0.023 0.119 0.150 0.041 0.003 0.166 0.369 0.396 0.125 0.018 0.039 0.162 0.192 0.055 0.004
10 0.024 0.122 0.153 0.042 0.003 0.400 0.610 0.626 0.209 0.046 0.092 0.266 0.299 0.091 0.010
12 0.026 0.127 0.161 0.044 0.003 0.385 0.596 0.617 0.203 0.044 0.156 0.364 0.389 0.123 0.017
Avg 0.024 0.121 0.152 0.042 0.003 0.299 0.512 0.532 0.174 0.034 0.079 0.232 0.262 0.079 0.009
Table 26: Full forecasting results for the Energy, Environment, and Health datasets using Autoformer, Informer, and Transformer as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models Autoformer Informer Transformer
(2021) (2021) (2017)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Energy Uni-modal 12 0.161 0.282 0.362 1.403 24.893 0.165 0.289 0.373 1.491 35.971 0.124 0.247 0.320 1.021 8.908
24 0.279 0.416 0.509 2.632 158.931 0.248 0.380 0.475 2.916 185.671 0.291 0.403 0.494 1.951 101.566
36 0.364 0.464 0.564 2.384 93.988 0.368 0.492 0.585 4.022 356.146 0.329 0.435 0.542 3.964 428.707
48 0.471 0.550 0.664 3.106 149.374 0.457 0.539 0.640 5.613 589.886 0.445 0.537 0.642 3.824 285.963
Avg 0.319 0.428 0.525 2.381 106.797 0.309 0.425 0.518 3.511 291.918 0.297 0.405 0.500 2.690 206.286
$M$-TSFLib 12 0.164 0.297 0.369 1.720 78.617 0.166 0.291 0.372 1.667 36.449 0.127 0.270 0.334 1.278 11.692
24 0.275 0.407 0.504 2.177 81.316 0.242 0.362 0.456 3.040 235.920 0.276 0.402 0.491 2.297 115.390
36 0.362 0.464 0.561 2.176 88.219 0.342 0.467 0.566 3.221 271.106 0.316 0.439 0.530 3.504 272.971
48 0.477 0.545 0.665 3.100 192.217 0.456 0.530 0.643 5.430 602.773 0.454 0.508 0.641 5.511 748.943
Avg 0.320 0.428 0.525 2.293 110.092 0.301 0.413 0.509 3.340 286.562 0.293 0.405 0.499 3.147 287.249
TaTS (ours) 12 0.153 0.291 0.365 1.490 33.553 0.145 0.271 0.347 1.118 11.872 0.126 0.252 0.324 1.036 12.192
24 0.276 0.404 0.488 2.040 73.651 0.242 0.361 0.463 2.310 184.196 0.265 0.381 0.469 1.706 53.962
36 0.360 0.467 0.555 2.167 80.434 0.301 0.415 0.514 3.658 379.160 0.301 0.419 0.515 2.804 209.452
48 0.466 0.559 0.659 3.431 178.191 0.448 0.538 0.644 4.344 378.638 0.423 0.527 0.620 5.149 416.124
Avg 0.314 0.430 0.517 2.282 91.457 0.284 0.396 0.492 2.857 238.466 0.279 0.395 0.482 2.674 172.933
Environment Uni-modal 48 0.503 0.544 0.670 1.953 141.941 0.414 0.481 0.602 2.190 216.476 0.412 0.483 0.600 2.190 214.543
96 0.594 0.605 0.743 2.015 135.052 0.467 0.520 0.644 2.454 290.858 0.458 0.511 0.639 2.608 340.210
192 0.607 0.601 0.762 2.163 205.175 0.469 0.519 0.644 2.535 305.217 0.492 0.535 0.677 2.711 404.734
336 0.690 0.643 0.823 2.624 307.413 0.487 0.529 0.693 2.726 355.026 0.479 0.517 0.688 2.824 389.196
Avg 0.599 0.598 0.749 2.189 197.395 0.459 0.512 0.646 2.476 291.894 0.460 0.511 0.651 2.583 337.171
$M$-TSFLib 48 0.442 0.492 0.626 2.008 164.960 0.419 0.473 0.604 1.997 163.772 0.420 0.471 0.606 2.027 163.407
96 0.456 0.505 0.645 1.930 154.161 0.426 0.477 0.616 2.074 173.217 0.426 0.485 0.615 2.380 252.997
192 0.461 0.505 0.663 1.995 181.553 0.425 0.481 0.634 2.410 256.710 0.418 0.479 0.630 2.356 252.249
336 0.447 0.506 0.666 1.994 172.427 0.427 0.489 0.651 2.283 225.380 0.435 0.487 0.657 2.551 282.514
Avg 0.452 0.502 0.650 1.982 168.275 0.424 0.480 0.626 2.191 204.770 0.425 0.481 0.627 2.329 237.792
TaTS (ours) 48 0.274 0.374 0.486 1.159 21.042 0.272 0.377 0.483 1.103 18.922 0.268 0.378 0.481 1.074 17.152
96 0.286 0.387 0.509 1.183 21.118 0.287 0.401 0.507 1.040 15.636 0.272 0.391 0.496 1.028 15.508
192 0.291 0.382 0.528 1.272 24.699 0.297 0.426 0.531 1.003 13.899 0.277 0.398 0.512 1.030 16.253
336 0.288 0.403 0.534 1.122 20.057 0.286 0.419 0.533 0.989 13.846 0.287 0.420 0.534 1.006 15.226
Avg 0.285 0.387 0.514 1.184 21.729 0.285 0.406 0.513 1.034 15.576 0.276 0.397 0.506 1.035 16.035
Health Uni-modal 12 1.389 0.895 1.121 4.898 593.771 1.173 0.743 0.990 4.065 593.314 1.143 0.741 0.998 3.233 133.066
24 2.328 1.191 1.476 6.117 859.160 1.215 0.766 1.042 3.728 389.282 1.480 0.786 1.096 2.435 59.228
36 1.953 1.005 1.348 4.614 392.542 1.315 0.773 1.101 3.541 390.034 1.450 0.772 1.132 2.766 136.630
48 2.179 1.064 1.410 4.865 379.673 1.408 0.809 1.137 3.445 337.635 1.439 0.806 1.137 2.834 125.016
Avg 1.962 1.039 1.339 5.123 556.286 1.278 0.773 1.067 3.695 427.566 1.378 0.776 1.091 2.817 113.485
$M$-TSFLib 12 1.331 0.864 1.112 4.427 504.600 0.997 0.673 0.928 3.024 221.350 0.953 0.686 0.920 3.309 198.713
24 1.454 0.868 1.138 3.681 181.137 1.209 0.740 1.046 3.372 257.237 1.315 0.762 1.071 2.833 137.728
36 1.574 0.906 1.216 4.051 234.643 1.282 0.760 1.092 3.111 213.531 1.276 0.763 1.070 2.906 137.944
48 1.616 0.911 1.235 3.884 201.678 1.371 0.785 1.137 3.146 234.342 1.330 0.780 1.109 2.769 125.761
Avg 1.494 0.887 1.175 4.011 280.514 1.215 0.740 1.051 3.163 231.615 1.218 0.748 1.042 2.954 150.036
TaTS (ours) 12 1.320 0.863 1.105 3.769 309.199 0.949 0.658 0.910 3.707 391.895 0.909 0.655 0.862 3.251 292.588
24 1.467 0.885 1.159 3.485 152.650 1.130 0.736 1.020 3.897 352.680 1.187 0.728 1.030 3.143 192.685
36 1.368 0.825 1.120 3.962 230.059 1.301 0.779 1.110 3.540 270.359 1.175 0.728 1.048 3.096 193.526
48 1.483 0.871 1.180 4.108 279.227 1.354 0.833 1.143 3.962 372.090 1.297 0.762 1.113 3.074 185.199
Avg 1.409 0.861 1.141 3.831 242.784 1.183 0.752 1.046 3.776 346.756 1.142 0.718 1.013 3.141 216.000
Table 27: Full forecasting results for the Security, Social Good, and Traffic datasets using Autoformer, Informer, and Transformer as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Models Autoformer Informer Transformer
(2021) (2021) (2017)
Method
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Security Uni-modal 6 115.544 5.268 10.749 2.437 185.108 133.168 6.626 11.540 1.591 19.155 126.925 6.203 11.266 2.547 120.160
8 113.355 4.967 10.647 2.888 298.107 134.383 6.804 11.592 1.236 6.780 132.905 6.641 11.528 1.209 14.174
10 114.407 4.984 10.696 2.164 153.069 131.382 6.631 11.462 1.483 26.411 128.820 6.443 11.350 1.539 32.792
12 117.801 5.253 10.854 2.692 326.651 128.201 6.430 11.323 1.408 26.498 136.753 7.039 11.694 1.168 4.854
Avg 115.277 5.118 10.736 2.545 240.734 131.784 6.623 11.479 1.429 19.711 131.351 6.582 11.460 1.616 42.995
$M$-TSFLib 6 108.841 4.861 10.433 3.004 246.068 127.216 6.197 11.279 2.326 99.099 127.277 6.233 11.282 2.592 135.752
8 113.693 5.031 10.663 2.799 265.988 128.914 6.402 11.354 1.660 40.618 126.589 6.227 11.251 2.010 84.511
10 112.545 5.098 10.609 1.885 120.031 128.775 6.434 11.348 1.532 38.764 127.986 6.368 11.313 1.520 32.213
12 110.699 4.902 10.521 2.177 199.450 130.907 6.626 11.441 1.443 33.523 132.033 6.685 11.491 1.217 11.881
Avg 111.445 4.973 10.556 2.466 207.884 128.953 6.415 11.356 1.740 53.001 128.471 6.378 11.334 1.835 66.089
TaTS (ours) 6 107.430 4.658 10.365 4.255 528.013 127.814 6.269 11.306 2.263 110.185 122.803 5.927 11.082 2.966 185.980
8 107.108 4.750 10.349 3.531 455.878 125.990 6.217 11.225 2.390 135.504 123.882 6.057 11.130 2.505 151.979
10 110.061 4.891 10.491 3.615 562.774 125.737 6.248 11.213 2.231 134.082 125.666 6.262 11.210 2.232 128.702
12 109.361 4.847 10.458 3.060 457.480 127.103 6.369 11.274 1.885 86.102 125.971 6.290 11.224 1.899 90.926
Avg 108.490 4.787 10.416 3.615 501.036 126.661 6.276 11.255 2.192 116.468 124.581 6.134 11.162 2.401 139.397
Social Good Uni-modal 6 1.234 0.684 0.973 1.910 69.347 0.773 0.425 0.668 0.872 13.761 0.848 0.427 0.684 1.208 27.858
8 1.236 0.664 0.956 1.998 80.017 0.871 0.449 0.708 0.933 11.089 0.877 0.460 0.719 1.520 56.020
10 1.332 0.708 1.026 2.116 93.499 0.910 0.498 0.755 0.905 6.360 0.976 0.541 0.783 1.725 63.227
12 1.312 0.749 1.046 2.231 75.925 0.926 0.644 0.818 1.028 7.619 0.938 0.508 0.750 1.389 28.551
Avg 1.278 0.701 1.000 2.064 79.697 0.870 0.504 0.737 0.934 9.707 0.910 0.484 0.734 1.460 43.914
$M$-TSFLib 6 1.200 0.664 0.954 1.890 66.499 0.772 0.414 0.666 0.726 8.039 0.784 0.422 0.673 0.938 18.507
8 1.175 0.589 0.891 1.935 84.135 0.829 0.405 0.673 0.746 6.869 0.845 0.426 0.669 1.056 19.505
10 1.268 0.689 1.002 1.756 44.172 0.876 0.466 0.724 0.855 6.248 0.913 0.486 0.731 1.252 25.176
12 1.272 0.739 1.032 2.048 56.789 0.879 0.541 0.765 0.888 7.772 0.882 0.510 0.744 1.021 13.844
Avg 1.229 0.670 0.970 1.907 62.899 0.839 0.457 0.707 0.804 7.232 0.856 0.461 0.704 1.067 19.258
TaTS (ours) 6 1.057 0.564 0.877 1.451 33.099 0.724 0.448 0.660 0.765 7.156 0.740 0.375 0.616 0.700 3.824
8 1.184 0.686 0.954 1.476 20.587 0.788 0.463 0.696 0.890 8.039 0.805 0.393 0.632 0.828 11.467
10 1.278 0.714 0.986 1.578 21.226 0.828 0.468 0.691 0.880 8.487 0.842 0.485 0.700 1.062 11.719
12 1.261 0.699 0.979 1.544 21.227 0.899 0.455 0.698 0.785 4.917 0.842 0.424 0.650 0.777 5.973
Avg 1.195 0.666 0.949 1.512 24.035 0.810 0.459 0.686 0.830 7.150 0.807 0.419 0.649 0.842 8.246
Traffic Uni-modal 6 0.201 0.295 0.416 0.316 0.542 0.197 0.352 0.434 0.303 0.328 0.203 0.344 0.438 0.299 0.307
8 0.205 0.296 0.419 0.299 0.440 0.202 0.358 0.440 0.320 0.369 0.224 0.364 0.460 0.302 0.285
10 0.200 0.297 0.413 0.310 0.479 0.196 0.351 0.430 0.296 0.314 0.209 0.353 0.445 0.298 0.295
12 0.241 0.303 0.448 0.350 0.653 0.213 0.358 0.448 0.326 0.397 0.200 0.323 0.426 0.316 0.416
Avg 0.212 0.298 0.424 0.319 0.528 0.202 0.355 0.438 0.311 0.352 0.209 0.346 0.442 0.304 0.326
$M$-TSFLib 6 0.195 0.248 0.390 0.247 0.360 0.166 0.297 0.387 0.285 0.391 0.164 0.291 0.384 0.271 0.340
8 0.204 0.274 0.401 0.277 0.404 0.169 0.306 0.393 0.285 0.372 0.172 0.305 0.395 0.274 0.318
10 0.207 0.274 0.409 0.268 0.407 0.169 0.305 0.391 0.282 0.358 0.169 0.297 0.391 0.283 0.393
12 0.241 0.292 0.420 0.343 0.795 0.183 0.290 0.392 0.292 0.414 0.181 0.288 0.392 0.287 0.389
Avg 0.212 0.272 0.405 0.284 0.492 0.172 0.299 0.391 0.286 0.384 0.171 0.295 0.390 0.279 0.360
TaTS (ours) 6 0.161 0.225 0.356 0.253 0.440 0.159 0.282 0.375 0.266 0.337 0.159 0.274 0.373 0.269 0.384
8 0.167 0.237 0.353 0.258 0.416 0.161 0.289 0.378 0.270 0.340 0.156 0.268 0.367 0.266 0.386
10 0.163 0.214 0.339 0.250 0.477 0.157 0.276 0.368 0.274 0.399 0.157 0.270 0.367 0.267 0.383
12 0.217 0.242 0.379 0.293 0.575 0.181 0.279 0.387 0.289 0.429 0.183 0.285 0.391 0.292 0.423
Avg 0.177 0.229 0.357 0.264 0.477 0.164 0.281 0.377 0.275 0.376 0.164 0.274 0.374 0.274 0.394
Table 28: Full forecasting results for various datasets using different time series modeling methods. Our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Avg: the average results across all prediction lengths.
Methods TaTS (ours) TaTS (ours) TaTS (ours) N-BEATS N-HiTS TCN ChatTime GPT4MTS
+ iTransformer + PatchTST + FiLM (2020) (2022) (2018) (2025) (2024)
Datasets
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Time-MMD (2024a) Agriculture 6 0.067 0.184 0.066 0.171 0.087 0.205 2.436 1.210 2.274 1.134 3.957 1.770 0.502 0.441 0.197 0.292
8 0.094 0.210 0.096 0.217 0.110 0.223 3.671 1.446 1.575 0.866 5.078 1.919 0.510 0.449 0.352 0.426
10 0.122 0.251 0.126 0.260 0.146 0.249 3.078 1.416 1.441 0.910 4.464 1.866 0.505 0.445 0.358 0.413
12 0.153 0.271 0.166 0.292 0.196 0.328 3.883 1.759 2.119 1.217 3.173 1.633 0.517 0.452 0.404 0.444
Avg 0.109 0.229 0.114 0.235 0.135 0.251 3.267 1.458 1.852 1.032 4.168 1.797 0.508 0.447 0.327 0.393
Climate 6 1.020 0.797 0.976 0.782 0.912 0.758 1.123 0.894 1.114 0.881 1.137 0.875 1.507 1.007 1.062 0.844
8 1.025 0.797 0.995 0.803 0.917 0.751 1.070 0.844 0.981 0.791 1.046 0.804 1.524 1.012 1.177 0.893
10 1.033 0.808 1.022 0.796 0.947 0.759 1.093 0.850 1.152 0.877 1.043 0.857 1.578 1.027 1.100 0.863
12 1.033 0.812 1.022 0.810 1.005 0.821 1.085 0.856 1.166 0.884 1.167 0.926 1.664 1.031 1.169 0.893
Avg 1.028 0.804 1.004 0.798 0.945 0.772 1.093 0.861 1.103 0.858 1.098 0.866 1.568 1.019 1.127 0.873
Economy 6 0.008 0.077 0.009 0.080 0.009 0.080 0.727 0.782 0.224 0.432 5.390 2.315 0.042 0.156 0.013 0.091
8 0.008 0.077 0.008 0.078 0.009 0.079 0.874 0.882 0.675 0.793 5.811 2.406 0.045 0.157 0.014 0.096
10 0.009 0.079 0.009 0.079 0.009 0.079 0.999 0.849 0.486 0.647 5.426 2.324 0.052 0.165 0.015 0.098
12 0.008 0.076 0.009 0.080 0.009 0.081 1.438 1.166 0.389 0.465 5.556 2.352 0.059 0.188 0.016 0.101
Avg 0.008 0.077 0.009 0.079 0.009 0.080 1.010 0.920 0.444 0.584 5.546 2.349 0.049 0.166 0.014 0.096
Energy 12 0.106 0.234 0.106 0.234 0.118 0.245 0.146 0.271 0.195 0.336 0.165 0.318 0.128 0.253 0.117 0.244
24 0.226 0.355 0.206 0.336 0.222 0.347 0.280 0.413 0.612 0.582 0.407 0.507 0.258 0.363 0.212 0.347
36 0.306 0.411 0.305 0.412 0.311 0.414 0.431 0.482 0.380 0.474 0.512 0.569 0.361 0.465 0.328 0.424
48 0.421 0.502 0.416 0.501 0.434 0.510 0.460 0.532 0.301 0.460 0.634 0.654 0.473 0.588 0.421 0.497
Avg 0.265 0.376 0.258 0.371 0.271 0.379 0.329 0.424 0.372 0.463 0.430 0.512 0.305 0.417 0.269 0.378
Environment 48 0.268 0.370 0.271 0.377 0.269 0.373 0.448 0.504 0.440 0.523 0.805 0.693 0.583 0.594 0.312 0.390
96 0.267 0.370 0.279 0.376 0.279 0.377 0.540 0.582 0.421 0.522 0.738 0.686 0.575 0.591 0.345 0.408
192 0.272 0.366 0.272 0.366 0.271 0.367 0.522 0.592 0.635 0.660 1.466 0.982 0.577 0.594 0.358 0.440
336 0.261 0.369 0.269 0.366 0.267 0.366 0.563 0.614 0.593 0.628 0.407 0.490 0.585 0.599 0.377 0.453
Avg 0.267 0.369 0.273 0.371 0.272 0.371 0.518 0.573 0.522 0.583 0.854 0.713 0.580 0.594 0.348 0.422
Health 12 0.939 0.649 0.990 0.659 1.211 0.794 1.572 0.838 1.723 0.866 2.484 1.027 1.482 0.802 1.157 0.704
24 1.251 0.712 1.288 0.764 1.414 0.827 1.680 0.981 1.621 0.905 2.072 1.070 1.645 0.956 1.743 0.904
36 1.489 0.781 1.397 0.780 1.497 0.851 1.834 0.980 1.687 0.908 1.557 0.859 1.732 0.937 1.950 0.938
48 1.581 0.834 1.456 0.808 1.562 0.878 1.556 0.952 1.635 0.911 1.639 0.922 1.813 0.942 2.217 0.957
Avg 1.315 0.744 1.283 0.753 1.421 0.838 1.660 0.938 1.666 0.898 1.938 0.970 1.668 0.909 1.766 0.875
Security 6 107.113 4.856 106.160 4.696 105.482 4.511 131.917 6.866 135.416 6.896 129.616 6.801 130.751 6.854 118.425 5.102
8 112.560 5.204 108.803 5.052 107.657 4.792 113.095 6.287 144.874 7.780 128.578 6.936 134.526 6.875 118.952 5.106
10 113.789 5.227 111.110 5.090 109.886 4.771 163.555 7.344 156.898 6.797 160.129 7.077 131.632 6.895 119.521 5.253
12 114.754 5.318 112.699 5.237 108.375 4.870 111.694 5.974 115.307 6.437 128.062 6.676 135.513 6.923 120.732 5.331
Avg 112.054 5.151 109.693 5.019 107.850 4.736 130.065 6.618 138.124 6.978 136.596 6.873 133.106 6.887 119.407 5.198
Social Good 6 0.942 0.398 0.923 0.436 0.992 0.588 1.752 0.654 1.446 0.626 1.535 0.917 1.213 0.608 1.214 0.485
8 0.967 0.433 0.900 0.461 1.001 0.550 0.952 0.551 1.033 0.531 1.129 0.935 1.252 0.621 1.422 0.560
10 0.994 0.463 0.996 0.461 1.065 0.583 1.116 0.627 1.112 0.594 1.208 0.996 1.278 0.667 1.264 0.572
12 1.045 0.514 1.069 0.501 1.358 0.782 1.445 0.827 1.498 0.846 1.385 0.996 1.313 0.712 1.757 0.622
Avg 0.987 0.452 0.972 0.465 1.104 0.626 1.316 0.665 1.272 0.649 1.314 0.961 1.264 0.652 1.414 0.559
Traffic 6 0.174 0.218 0.155 0.204 0.157 0.228 0.352 0.409 0.295 0.372 0.779 0.775 0.369 0.435 0.185 0.227
8 0.177 0.213 0.162 0.210 0.162 0.260 0.329 0.467 0.279 0.385 0.674 0.704 0.361 0.432 0.190 0.240
10 0.186 0.225 0.167 0.214 0.169 0.269 0.342 0.474 0.281 0.404 0.702 0.758 0.363 0.427 0.189 0.246
12 0.213 0.212 0.361 0.241 0.215 0.236 0.364 0.505 0.217 0.368 0.675 0.739 0.359 0.422 0.219 0.272
Avg 0.187 0.217 0.172 0.209 0.176 0.248 0.347 0.464 0.268 0.382 0.708 0.744 0.363 0.429 0.195 0.246
FNSPID (2024) Delta Airlines (DAL) 6 0.064 0.161 0.059 0.160 0.065 0.167 0.253 0.417 0.191 0.342 0.247 0.422 0.068 0.169 0.069 0.179
12 0.110 0.233 0.116 0.224 0.126 0.235 0.320 0.470 0.260 0.416 0.341 0.509 0.128 0.235 0.118 0.223
Avg 0.087 0.197 0.086 0.192 0.095 0.201 0.286 0.444 0.226 0.379 0.294 0.466 0.098 0.202 0.093 0.201
IBM (IBM) 6 0.324 0.401 0.341 0.401 0.642 0.534 0.820 0.667 0.913 0.675 2.001 1.150 0.396 0.453 0.392 0.424
12 0.804 0.602 0.758 0.580 1.140 0.856 1.390 0.887 1.517 0.937 1.870 1.096 0.807 0.619 0.886 0.627
Avg 0.564 0.501 0.550 0.490 0.891 0.695 1.105 0.777 1.215 0.806 1.936 1.123 0.602 0.536 0.639 0.525
JPMorgan Chase (JPM) 6 1.173 0.817 1.275 0.819 1.606 0.912 1.788 1.063 2.217 1.082 3.971 1.785 1.332 0.843 1.644 0.922
12 2.214 1.124 2.470 1.161 3.421 1.281 3.050 1.287 4.635 1.383 3.557 1.637 2.742 1.244 2.622 1.323
Avg 1.693 0.970 1.872 0.990 2.513 1.096 2.419 1.175 3.426 1.232 3.764 1.711 2.037 1.043 2.133 1.122
NVIDIA (NVDA) 6 0.032 0.123 0.041 0.151 0.037 0.172 0.257 0.424 0.093 0.224 0.431 0.560 0.038 0.139 0.043 0.149
12 0.054 0.159 0.055 0.162 0.063 0.177 0.288 0.445 0.151 0.301 0.483 0.587 0.067 0.183 0.066 0.170
Avg 0.043 0.141 0.048 0.156 0.050 0.174 0.272 0.434 0.122 0.262 0.457 0.574 0.053 0.161 0.054 0.159
Pfizer (PFE) 6 0.240 0.355 0.255 0.363 0.329 0.419 0.767 0.623 0.662 0.608 0.601 0.557 0.354 0.451 0.283 0.389
12 0.412 0.476 0.439 0.482 0.567 0.535 0.586 0.520 0.987 0.718 0.654 0.561 0.462 0.503 0.456 0.490
Avg 0.326 0.416 0.347 0.422 0.448 0.477 0.676 0.572 0.824 0.663 0.628 0.559 0.408 0.477 0.369 0.439
Tesla (TSLA) 6 0.101 0.241 0.071 0.204 0.098 0.235 0.166 0.306 0.262 0.452 3.784 1.929 0.123 0.256 0.166 0.315
12 0.183 0.321 0.245 0.390 0.122 0.253 0.210 0.358 0.247 0.381 3.168 1.773 0.192 0.343 0.196 0.292
Avg 0.142 0.281 0.158 0.297 0.110 0.244 0.188 0.332 0.254 0.416 3.476 1.851 0.158 0.300 0.181 0.303
FNF (2024b) Bitcoin Price 6 1.760 0.916 1.552 0.858 1.751 0.899 123.135 9.413 105.978 9.806 43.935 5.848 2.525 1.085 1.808 0.894
8 2.147 1.006 1.998 0.964 2.458 1.080 126.985 9.703 163.110 10.690 53.318 6.412 3.352 1.163 2.521 1.050
10 3.007 1.212 2.621 1.128 3.399 1.287 157.388 11.113 179.247 10.759 65.162 7.253 3.853 1.486 3.014 1.208
12 3.521 1.313 3.185 1.229 3.491 1.304 159.445 11.192 187.741 10.884 68.887 7.464 4.631 1.641 3.542 1.456
Avg 2.609 1.112 2.339 1.045 2.775 1.142 141.738 10.355 159.019 10.535 57.826 6.744 3.590 1.344 2.721 1.152
Web Traffic 6 15.106 2.504 17.348 2.755 17.378 2.755 21.596 3.003 21.993 2.912 16.625 2.157 19.265 2.869 18.106 2.832
8 18.474 2.740 17.715 2.625 16.11 2.600 19.330 2.754 21.001 2.828 20.144 2.826 21.724 2.924 18.421 2.745
10 18.313 2.783 18.008 2.643 16.23 2.516 20.866 2.901 24.046 2.930 21.976 2.947 20.164 2.875 19.902 2.792
12 19.084 2.823 18.783 2.590 22.721 3.217 27.141 3.409 26.011 3.374 27.759 3.439 21.841 2.969 20.611 2.942
Avg 17.744 2.712 17.964 2.653 18.110 2.772 22.233 3.017 23.263 3.011 21.626 2.842 20.748 2.909 19.260 2.827
Electricity Demand 6 0.262 0.379 0.257 0.374 0.260 0.375 0.385 0.470 0.338 0.443 0.436 0.506 0.510 0.544 0.271 0.388
8 0.274 0.391 0.257 0.370 0.244 0.360 0.419 0.497 0.426 0.504 0.501 0.565 0.524 0.562 0.284 0.369
10 0.288 0.405 0.173 0.266 0.252 0.366 0.407 0.494 0.335 0.454 0.558 0.582 0.536 0.571 0.292 0.398
12 0.295 0.408 0.327 0.424 0.310 0.408 0.451 0.522 0.355 0.471 0.567 0.601 0.558 0.588 0.337 0.428
Avg 0.280 0.396 0.254 0.358 0.266 0.377 0.416 0.496 0.364 0.468 0.516 0.564 0.532 0.566 0.296 0.395
E.4 Full Imputation Results
Table 29: Full imputation results for the Climate, Economy, and Traffic datasets using PatchTST, DLinear and FiLM as time series models. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework seamlessly enhances existing time series models to effectively handle time series with concurrent texts. Promotion: the improvement of the best baseline.
Models PatchTST DLinear FiLM
(2023) (2023) (2022a)
Metric
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
MSE
MAE
RMSE
MAPE
MSPE
Climate
Uni-modal
1.111 0.846 1.052 3.898 442.010 0.969 0.801 0.983 1.881 59.681 1.123 0.829 1.059 2.647 128.421
$M$-TSFLib
1.010 0.821 1.002 2.156 81.983 0.963 0.802 0.980 1.394 18.580 1.130 0.833 1.061 1.949 51.811
TaTS (ours)
0.878 0.720 0.937 1.573 35.877 0.912 0.757 0.951 1.701 25.242 0.820 0.718 0.902 1.478 30.622
Promotion
13.1% 12.3% 6.5% 27.0% 56.2% 5.3% 5.5% 3.0% -22.0% -35.9% 27.0% 13.4% 14.8% 24.2% 40.9%
Economy
Uni-modal
0.029 0.138 0.170 0.051 0.004 0.057 0.190 0.239 0.068 0.007 0.077 0.209 0.277 0.076 0.010
$M$-TSFLib
0.026 0.137 0.161 0.049 0.003 0.061 0.196 0.247 0.069 0.007 0.075 0.203 0.271 0.072 0.009
TaTS (ours)
0.017 0.107 0.128 0.038 0.002 0.045 0.171 0.210 0.061 0.006 0.054 0.168 0.232 0.061 0.007
Promotion
34.6% 21.9% 20.5% 22.4% 33.3% 21.1% 10.0% 12.1% 10.3% 14.3% 28.0% 17.2% 14.4% 15.3% 22.2%
Traffic
Uni-modal
0.210 0.339 0.444 0.358 0.600 0.245 0.417 0.489 0.430 0.720 0.175 0.311 0.409 0.343 0.508
$M$-TSFLib
0.189 0.341 0.428 0.391 0.690 0.179 0.335 0.419 0.362 0.630 0.169 0.288 0.396 0.316 0.545
TaTS (ours)
0.131 0.248 0.331 0.312 0.647 0.134 0.297 0.352 0.266 0.270 0.137 0.242 0.354 0.281 0.388
Promotion
30.7% 26.8% 22.7% 12.8% -7.8% 25.1% 11.3% 16.0% 26.5% 57.1% 18.9% 16.0% 10.6% 11.1% 23.6%
E.5 Full Hyperparameter Study Results of Learning Rate
(a) Agriculture
(b) Climate
(c) Economy
(d) Energy
(e) Environment
(f) Health
(g) Security
(h) Social Good
(i) Traffic
Figure 11: Parameter study on the learning rate. We evaluate the impact of varying the learning rate in $\{0.00005,0.0001,0.00015,0.0002,0.00025,0.0003\}$ by reporting the mean squared error (MSE) of our TaTS framework across datasets. The results demonstrate that TaTS maintains stable performance across different learning rate choices.
E.6 Full Hyperparameter Study Results of Text Embedding Dimension
(a) Agriculture
(b) Climate
(c) Economy
(d) Energy
(e) Environment
(f) Health
(g) Security
(h) Social Good
(i) Traffic
Figure 12: Parameter study on the projection dimension of paired texts. We vary the text projection dimension in $\{6,12,18,24\}$ and report the mean squared error (MSE) of our TaTS framework across datasets. The results indicate that TaTS maintains robust performance across different choices of text projection dimensions.
E.7 Full Ablation Study Results Using Different Text Encoders
We conduct experiments to evaluate the performance of our TaTS with multiple language encoders. Specifically, we evaluation TaTS with BERT-110M [^2] https://huggingface.co/google-bert/bert-base-uncased, GPT2-1.5B [^3] https://huggingface.co/openai-community/gpt2 and LLaMA2-7B [^4] https://huggingface.co/meta-llama/Llama-2-7b as the language encoders. The results, presented in Figure 13, demonstrate that TaTS remains robust across different text encoders and consistently outperforms the baselines.
(a) Agriculture
(b) Climate
(c) Economy
(d) Energy
(e) Environment
(f) Health
(g) Security
(h) Social Good
(i) Traffic
Figure 13: Performance comparison of different text encoders within the TaTS framework. Specifically, we evaluate BERT-110M, GPT2-1.5B, and LLaMA2-7B across multiple datasets using PatchTST (transformer-based model), DLinear (linear-based model), and FiLM (frequency-based model). TaTS maintains relatively stable performance across various models and datasets.
In the main experiments, the text embeddings are obtained by a mean pooling over token embeddings. There are more advanced embedding techniques. To this end, we conducted an additional experiment using LLaMA-3.2-1B to generate sentence-level embeddings, while keeping all other settings the same. The results are shown below in Table 30.
Table 30: Mean Square Error of TaTS forecasting with different text embedding methods.
Text Embedding Method Agriculture Climate Economy Security Social Good Traffic
GPT-2 average pooling (prediction length 6) 0.067 1.020 0.008 107.113 0.942 0.174
LLaMA-3.2-1B sentence embeddings (prediction length 6) 0.067 1.015 0.008 106.625 0.935 0.179
GPT-2 average pooling (prediction length 12) 0.153 1.033 0.008 114.754 1.045 0.213
LLaMA-3.2-1B sentence embeddings (prediction length 12) 0.152 1.021 0.008 114.219 1.025 0.209
E.8 Full Efficiency Results: Computational Overhead vs. Performance Gain Trade-offs
We conduct experiments to analyze the efficiency of TaTS, with results presented in Figure 14. Each subfigure visualizes the training time per epoch and the forecasting mean squared error (MSE) for different time series models, represented as transparent colored scatter points. The average performance is computed and marked with cross markers: the green cross represents the average performance of TaTS, while the red and blue crosses indicate the average performance of the baseline models. As TaTS introduces a lightweight MLP and augments the original time series with auxiliary variables projected from paired texts, it incurs a slight computational overhead, with an average increase of $\sim 8\%$. Yet this trade-off results in a $\sim 14\%$ average improvement of forecasting MSE.
(a) Agriculture
(b) Climate
(c) Economy
(d) Energy
(e) Environment
(f) Health
(g) Security
(h) Social Good
(i) Traffic
Figure 14: Efficiency and forecasting performance comparison of uni-modal time-series modeling, $M$-TSFLib, and our TaTS framework. While TaTS incurs a slight increase in training time due to the augmentation of auxiliary variables projected from paired texts, it significantly enhances forecasting accuracy, achieving a lower MSE.